### v1

# Accelerating FRB Search: Dataset and Methods

Abstract
Fast Radio Burst (FRB) is an extremely energetic cosmic phenomenon of short duration. Discovered only recently and with yet unknown origin, FRBs have already started to play a significant role in studying the distribution and evolution of matter in the universe. FRBs can only be observed through radio telescopes, which produce petabytes of data, rendering the search for FRB a challenging task. Traditional techniques are computationally expensive, time-consuming, and generally biased against weak signals. Various machine learning algorithms have been developed and employed, which all require substantial data sets. We here introduce the FAST dataset for Fast Radio bursts EXploration (FAST-FREX), built upon the observations obtained by the Five-hundred-meter Aperture Spherical radio Telescope (FAST). Our dataset comprises 600 positive samples of observed FRB signals from three sources and 1000 negative samples of noise and Radio Frequency Interference (RFI). Furthermore, we provide a machine learning algorithm, Radio Single-Pulse Detection Algorithm Based on Visual Morphological Features (RaSPDAM), with significant improvements in efficiency and accuracy for FRB search. We also employed the benchmark comparison between conventional single-pulse search softwares, namely PRESTO and Heimdall, and RaSPDAM. Future machine learning algorithms can use this as a reference point to measure their performance and help the potential improvements.
Fast radio bursts(2008) — Astronomy databases(83) — Astronomy software(1855)
1 Introduction
Fast Radio Burst (FRB) is a rapidly developing field in astrophysics that researchers extensively explore regarding its origin, emission mechanisms, and applications. FRBs are characterized by their short duration, high-energy release, and widespread distribution in the universe, which is crucial in studying the distribution and evolution of matter in the cosmos. Despite the discovery of about a few hundred FRBs (Li et al., 2021; Niu et al., 2022a; Feng et al., 2022; Zhang et al., 2022), their origin and physical mechanisms are still a mystery. The discovery of more FRBs is expected to bring revolutionary breakthroughs in physics and astronomy.
Various open-source software packages have been developed for searching single-pulse signals to discover FRBs, long-period pulsars, and other radio transient sources. PRESTO (Ransom, 2011) and SIGPROC (Lorimer, 2011) are two software packages designed to standardize the initial analysis of various fast-sampled pulsar data types and have aided in the discovery of numerous pulsar radio sources. Heimdall  [^1] https://sourceforge.net/p/heimdall-astro/wiki/Home/, a single-pulse search algorithm based on Graphical Processing Units (GPUs), can enhance data processing efficiency to a certain degree. These software packages’ pipelines usually employ computation-intensive dispersion removal algorithms on search data, producing several time sequences with dispersion effects removed, which might lead to significant computational time. On the other hand, Radio Frequency Interference (RFI) and instrument noise make detecting weaker pulse signals challenging.
In recent years, the field of astronomy has extensively applied machine learning methods, providing novel approaches and more efficient practical methods for detecting FRBs. Based on deep learning classifiers, these methods can effectively reduce the substantial number of false-positive single-pulse candidates, mainly due to RFI (Agarwal et al., 2020b, a; Connor & van Leeuwen, 2018; Chen et al., 2023). The above methods rely on FRB candidates generated from single-pulse search algorithms that use de-dispersion. The current FRB detector directly analyses raw observational data (Liu et al., 2022). These algorithms can process search data in real-time and detect weak signals that current search technologies might have missed. However, training machine learning models requires a substantial amount of data samples. Consequently, supplying a FRB dataset for the machine learning algorithm research is an impending issue that needs to be resolved urgently.
Astronomical telescopes have gathered extensive observational data that aids in the search for FRBs. The Chinese Five-hundred-meter Aperture Spherical radio Telescope (FAST) (Li et al., 2018; Jiang et al., 2019), which is the world’s biggest single-aperture telescope with the highest sensitivity, can detect weak signals that other radio telescopes cannot detect due to its precise surface reflection control and the favorable polarization characteristics of its FAST-19 beam receiver. The research team utilized FAST to detect 1652 outbursts in about a 50-day interval through the Commensal Radio Astronomy FAST Survey (CRAFTS), obtaining the most extensive sample of FRB outbursts to date and revealing the complete spectrum and bimodal structure of the burst rate for the first time (Li et al., 2021).
This paper proposes the FAST dataset for Fast Radio bursts EXploration (FAST-FREX). FAST-FREX was derived from the authentic observation data obtained from FAST, containing FRB signals from multiple sources. The FRB sources include FRB20121102(Li et al., 2021), FRB20180301(Laha et al., 2022) and FRB20201124(Zhang et al., 2022; Niu et al., 2022b). FAST-FREX aims to facilitate the advancement of machine learning algorithms for FRB search, enabling the exploration of more FRB events. We introduce a machine learning algorithm, Radio Single-Pulse Detection Algorithm Based on Visual Morphological Features (RaSPDAM), showcasing substantial improvements in efficiency and accuracy on FRB search. Furthermore, we present the benchmark results from two conventional softwares and RaSPDAM, serving as baselines for evaluating the performance of machine learning algorithms in future research.
The paper is structured as follows: Section 2 outlines the present state of dataset construction in the astronomy field. Section 3 details the methods and parameter specifications employed to create the dataset. Section 4 comprehensively explains the machine learning algorithm RaSPDAM and introduces two conventional search software, PRESTO and Heimdall. The benchmark results of these methods are presented in Section 5 as the baselines. Section 6 introduces the application of the algorithm RaSPDAM and the pulsars discovered by the algorithm. Section 7 highlights the limitations of the dataset and potential challenges. Finally, Section 8 presents the conclusion and outlines future directions.
2 Related Work
At present, astronomical data is primarily obtained from multiple telescopes that record data according to their specific standards, resulting in a lack of uniformity. Additionally, observational targets and recording patterns differ based on the research field. Some open-source datasets are available for pulsar, and FRB searches to address this issue.
The HTRU2 dataset (Lyon et al., 2016), which stems from The High Time Resolution Universe Pulsar Survey project (Keith et al., 2010), comprises a significant number of pulsar candidates, totaling 17,898 samples. The sample set consists of 1,639 authentic pulse samples and 16,259 negative samples caused by RFI/noise. However, it only provides parameter files rather than detailed observational data, limiting the available information.
Although the LOTAAS1 dataset (Sanidas et al., 2019) comprises 66 pulsars and 4987 non-pulsar candidates discovered during the project, it is currently not publicly available.
The SPARKESX datasets (Single-dish PARKES data sets for finding the uneXpected) (Yong et al., 2022) intend to identify pulsars, FRBs, and other signals while supporting the development of new search algorithms. It contains three mock surveys from the Parkes ”Murriyang” radio telescope, providing vast sample datasets, including authentic and simulated high-time resolution observations. However, the positive samples in the datasets are simulated pulse signals of different types rather than genuine observed FRBs. Therefore, it may be less helpful in conducting specific FRB searches.
The Blinkverse (Xu et al., 2023) database  [^2] https://blinkverse.alkaidos.cn/ collects fast radio bursts released by various observatories, including FAST, CHIME, GBT, and Arecibo, and provides thousands of previously unavailable dynamic spectra from FAST. This comprehensive database features detailed pulse attributes, dynamic spectrograms, and source information, but does not provide detailed observational data.
Currently, based on the substantial volume of observational data, researchers have begun integrating machine learning algorithms into FRB exploration. The high-quality machine learning models perform real-time, efficient searches. For instance, employing a convolutional neural network to FRB121102 integrates neural network detection and dispersion removal verification, resulting in better sensitivity, lower false detection rate, and faster processing speed than conventional algorithms. The convolutional neural network detected new pulses in data obtained on August 26, 2017, where 21 bursts had been previously identified (Zhang et al., 2018). Moreover, Bo Han Chen and colleagues presented a technique for detecting repeating FRBs based on Uniform Manifold Approximation and Projection (UMAP). They discovered that unsupervised UMAP classification had a 95% completeness rate for identifying repeating FRBs and identified 188 potential sources of FRB repeaters from 474 non-repeating sources (Chen et al., 2022). Thus, creating datasets for advancing research in machine learning algorithms is paramount.
In response to the issues described above, we will offer an open-access dataset FAST-FREX of FRB signals obtained from observations made by FAST. The dataset contains authentic FRB signal samples instead of simulated ones. Since the FRB signal samples originate entirely from the observation data from FAST, they can more precisely represent real-world scenarios.
3 Dataset
3.1 Data Collection
We created our dataset by using the observational data of FRB20121102(Li et al., 2021), FRB20180301(Laha et al., 2022) and FRB20201124(Zhang et al., 2022; Niu et al., 2022b). The FAST continuous monitoring of FRB20121102 commenced in August 2019, leading to the detection of 1,652 independent burst events between August 29th and October 29th within 59.5 hours. Meanwhile, FAST observed FRB20180301 on 2021 March 4th, 9th, and 19th and detected five bright radio bursts. More than 800 bursts were detected from FRB20201124 by FAST between 2021 September 25th and 28th. The dataset comprises 470 signals from FRB20121102, 5 signals from FRB20180301, and 125 signals from FRB20201124.
These observations gathered data using 4096 frequency channels over 1.05 GHz to 1.45 GHz, with 0.122 MHz frequency resolution. These channels recorded four polarization signals. FRB20121102 has a 98.304 $\mu$s sampling rate, while others have a 49.152 $\mu$s sampling rate. The raw observation data was stored in FITS  [^3] https://fits.gsfc.nasa.gov/ format and divided into blocks of 128 or 256 time samples. The samples were recorded in consecutive lines (or sub-integrations) in a file with 1024 samples per sub-integration.
Our objective is to develop and compare algorithms that can identify FRB events. Therefore, we selected 600 bursts to create positive data samples and ensured that each positive sample contained only one FRB event. In order to preserve the integrity of FRB events, we set the cropping duration of each positive sample file to about 6 seconds. We also ensure that the events appear randomly within the observation time covered by the file rather than setting a fixed Time of Arrival (ToA). This approach simulates FRB detection more realistically and enhances the dataset’s diversity, which is beneficial for algorithm research. We reduced the data size by averaging the first two polarization signals while maintaining the signal’s integrity and mitigating the noise’s effect to some extent. After eliminating the records of authentic events, we extracted 1000 negative samples of RFI and noise with the same duration as the positive samples.
3.2 Dataset Features
The dataset comprises two file types: sample files and parameter files. The sample files, stored in FITS format, contain pre-cropped observation data. Among them are 600 positive sample files containing FRB signals and 1000 negative sample files containing RFI and noise. The FRB20121102 sample file has a time sampling point of 60 * 1024, while other sources’ sample files have 120 * 1024. Moreover, the sample file’s number of polarization channels is reduced to one, which differs from the original data. The size of each file is approximately 244 MB or 488 MB, depending on its time sampling rate. Meanwhile, parameter files are stored in CSV format to record various FRB parameters for each positive sample file (see Table 1).
Table 1: Parameters of Dataset
Name Abbr Unit
Modified Julian Day MJD -
Time of Arrival ToA s
Dispersion Measure DM pc $cm^{-3}$
Each positive sample file contains only one FRB signal and its parameters are recorded in a corresponding parameter file specific to a fixed FRB source. In constrast, negative sample files do not have a corresponding parameter file.
[…]
$t_{2}-t_{1}=D\times(\frac{1}{v_{1}^{2}}-\frac{1}{v_{2}^{2}})\times DM$ (1)
3.2.1 Authentic Single Pulses
From bursts of multiple FRB sources, we carefully chose 600 signals, which have a wide-spread distribution across four dimensions: (1) Full Width at Half Maximum (FWHM), (2) bandwidth, (3) peak flux density ($S_{\rm{peak}}$), and (4) fluence (F), to form the positive samples. This approach ensures the diversity of FRB signals in our dataset. Figure 1 depicts the distribution of these four parameters.
The FWHM is a parameter commonly used to describe the width of the burst profile. It is given by the distance between points on the spectral curve at which the intensity reaches half its peak. The burst is distributed over a frequency interval, and the range of this frequency interval is called bandwidth. The spectral flux density is a quantity that indicates the observed strength of an astronomical source. It is a measure of the strength of a radio signal received from a discrete source, which is measured in watts per square meter per hertz ($\rm{W~{}m^{-2}~{}Hz^{-1}}$). Peak flux density $S_{\rm{peak}}$ is the maximum flux intensity of the pulse profile. The unit of flux density is the jansky (Jy), and 1 Jy equals $\rm{10^{-26}~{}W~{}m^{-2}~{}Hz^{-1}}$. The fluence F of a burst is computed by integrating the burst profile concerning time and bandwidth, which is used to describe the total energy of the burst. The unit of F is Jy ms.
Figure 1: Distribution of feature parameters for FRBs
The DM range of these signals is limited, lying between 411.51 and 570.0 pc $cm^{-3}$. Pulse widths are varied over a range of 0.34 to 78.52 ms. The minimum $S_{\rm{peak}}$ is 0.00169 Jy, while the maximum is 0.84239 Jy. Moreover, the bandwidth is between 16 and 500 MHz, and the F is between 0.0016 and 4.5616 Jy ms. These data are collected from the cited articles.
In Figure 2, we present the frequency-time domain representation of FRB signals, varying pulse width of positive samples. Pulse width indicates the duration of the FRB signals on the time axis. Figure 3 displays the frequency-time domain of FRB signals, with varying signal flux densities of positive samples. This visualization illustrates that signals with higher flux density are more prominently observed than the background noise.
Figure 2: Examples of positive samples with different pulse widths
Figure 3: Examples of positive samples with different flux densities
3.2.2 Radio Frequency Interference
Based on FAST observations, we can categorize RFI in FAST data into three types: (1)narrow-band RFI,(2)1 MHz-wide RFI, and (3)fixed frequency RFI with broader bandwidth (Jiang et al., 2020). Narrow-band RFI can have various origins, such as interference from electronic devices or local effects of the telescope. The 1 MHz wide RFI is caused by the standing wave, which looks like a regular sinusoidal wave or the single bump in FAST spectra. The fixed frequency RFI is due to satellite or civil aviation from the sky. Several examples of negative samples are shown in Figure 4, where the bright yellow patches are predominantly fixed frequency RFI (type 3). In contrast, the first two types of RFI are challenging to discern visually. We selected RFI and noise randomly from the original observation files instead of simulated RFI injection. Similar to the positive samples, these negative samples possess diversity and represent real-world scenarios, as depicted in Figure 4.
Figure 4: Examples of negative samples
4 Methods
4.1 RaSPDAM
The RaSPDAM algorithm aims to identify FRB/PSR(PulSaR) signals using a computer vision-based approach, demonstrating significant enhancements in both efficiency and accuracy.
In conventional approaches, the initial step of searching for FRBs involves de-dispersion, a process that requires significant time and computational resources (CPU/GPU cores) to perform the Fast Fourier Transform (FFT) operation. As a result, the process of searching for FRBs has become computationally intensive and time-consuming. In the context of RaSPDAM, observational data is processed as conventional images. The signal sequence is segmented and converted into an image format with dimensions of $512\times 512$ pixels. The image serves as the input for a semantic segmentation model.
4.1.1 Signal Preprocessing & Enhancement
The initial step is to convert the original signal sequence into standard images without de-dispersion. The signal sequence is divided into smaller slices within a 2-second width sliding window and a 1-second overlap. To standardize the model input, we resized the image slices to dimensions of $512\times 512$.
The second step in preparing for segmentation is image enhancement. Typically, the FRB signal appears as a curve overlapped with noise in a sequence of signals. Initially, we use convolution to enhance the characteristics of the curve. In image processing, convolution transforms an image by applying a kernel to each pixel and its local neighbors across the entire image. For this case, as shown in Figure 5, we utilize multiple convolution kernels generated from fixed curve slopes. After convolving the image slices with multiple convolution kernels, we apply morphological dilation to further enhance the signal. Finally, to maximize the utilization of image features, the enhanced images (post-convolution and after morphological dilation) and the original image have been combined into an RGB image, which will be used as the model input. Figure 5 illustrates an example of this process.
Figure 5: Signal Image Preprocessing Procedure
4.1.2 Model Training
The aim of FRB detection resembles semantic segmentation, also known as pixel-based classification. It requires a method to differentiate the FRB signal component from the background, where noise might interfere with segmentation outcomes. In this case, we choose the UNet as the model implementation due to several key advantages:
  1. 1.
    High-performance: UNet is recognized for generating accurate segmentation maps, particularly when handling high-resolution images or datasets with numerous classes.
  2. 2.
    Efficiency: UNet incorporates high-level and low-level features from the input image depending on skip connections. It enhances the model’s efficiency in utilizing training data, improving overall performance.
Unet is a Convolutional Neural Network specialized in classification tasks, localizing specific areas by classifying individual pixels. The architecture resembles a U-shape, with the left side part known as the encoder or contracting path and the right side as the decoder or expansion path.
Training the network requires a significant amount of FRB signal images. However, there is a lack of available training data in the field of astronomy. Our proposed alternative involves generating simulated FRB signals. As previously mentioned, we enhance the signal sequence slice using convolution kernels. To generate simulated signals, we apply a similar process by using random curve slopes and Gaussian noises. With the known slope of the FRB curve, it is possible to accurately determine its position and create mask images for training purposes.
4.1.3 Candidate Identification
Figure 6: Candidate Identification
The network segmentation process produces segmentation maps for the purpose of identifying potential candidates for FRB signals. However, it may still contain noise. In response to this issue, we have implemented multiple procedures for candidate identification.
Figure 6 shows that we utilize the ”regionprops” function for the analysis of connected areas. Each identified candidate is quantified as x, while the area of the region box and the filled area are represented as area_bbox and area_filled. Then, we use thresholds t_a. If t_a is smaller than the ratio of area_filled to area_bbox, it is flagged as noise, as indicated in Equation 2.
$F(x)=\left\{\begin{array}[]{l}\ \frac{x.area\_filled}{x.area\_bbox}<t_{a}\\ \ \frac{x.x_{2}-x.x_{1}}{window\_width}>t_{b}\\ \ \frac{x.y_{2}-x.y_{1}}{window\_height}>t_{c}\end{array}\right.$ (2)
Additionally, we calculate the candidate’s projection on both the x-axis and y-axis. Candidates exceeding both x-axis and y-axis projections above thresholds that denoted as t_b and t_c, as potential FRB signals.
4.2 Conventional Methods
In addition to the machine learning algorithm RaSPDAM described above, we established baseline models using conventional single-pulse search software, including PRESTO and Heimdall. The search process for astronomical signals contains RFI removal, de-dispersion, and matched filtering and generates single pulse candidates. Interstellar media can interfere with the arrival time of high-frequency and low-frequency electromagnetic waves to radio telescopes. This phenomenon results in energy dispersion of astronomical signals, broadening of pulse profiles, decreased signal-to-noise ratio, or even disappearance of the pulse signal. To overcome these effects, PRESTO and Heimdall process the original signal using the principle of non-coherent de-dispersion before conducting single-pulse searching. This step compensates for different frequency-time delays, eliminating the effects of frequency-related delays and improving the signal-to-noise ratio.
4.2.1 PRESTO
PRESTO has many options to optimize the FRB search tasks. We choose a simple set of parameters as representative PRESTO output. The search pipeline includes RFI removal, de-dispersion, and single-pulse search. We identified and marked narrowband and short-duration broadband interference in the data as RFIs, recording it in ”.mask” files to facilitate removal. Due to the range of dispersion in the dataset, we set the DM grid range from 350 pc $cm^{-3}$ to 650 pc $cm^{-3}$ and the DM grid density at 1. The de-dispersion task with the above DM grid was applied to the dataset, leading to a set of de-dispersed time series. A single-pulse search was conducted for each DM value to detect pulses surpassing a Signal-to-Noise Ratio (SNR) threshold of 3.0, which were then cataloged in a candidate list.
4.2.2 Heimdall
Heimdall is a GPU-accelerated transient detection pipeline. It computes the search step based on smearing calculations for each DM value to balance computational requirements more evenly across the entire DM range, resulting in better dispersion removal at higher dispersion measures. Heimdall pipeline includes de-dispersion, baseline removal, normalization, matched filtering, re-normalization, and SNR threshold filtering. In the matched filtering step, a DM range of 350 pc $cm^{-3}$ 650 pc $cm^{-3}$ was chosen as same as PRESTO.
5 Benchmarks and Results
5.1 Baseline Scoring Metrics
We aim to improve the efficiency and accuracy of identifying positive samples containing FRB signals. Consequently, our approach yields metrics, including recall and precision rates. The True Positives (TPs) and True Negatives (TNs) mean the correctly predicted positive candidates and negative samples. Meanwhile, False Positives (FPs) and False Negatives (FNs) reflect the erroneously predicted positive candidates and negative samples. We define that detecting FRB signal candidates in positive samples with a ToA error $\leq$ 0.2s and a DM error $\leq$ 50 pc $cm^{-3}$ are correctly predicted TPs. Candidates in positive samples that do not meet the specified error thresholds and candidates found in negative samples are FPs. Consequently, the number of FPs may exceed the number of sample files, as each sample file could yield multiple candidates. If no FRB signal candidates appear in negative samples, they are correctly predicted as negative samples (TNs). If no candidates are found in positive samples, it is treated as FNs.
For the RaSPDAM algorithm, considering its image identification method, the DM value cannot be measured. Therefore, the determination of TP relies solely on ToA. Additionally, it cannot precisely calculate the ToA. Hence, if the ToA of a candidate falls within the signal’s entire period or has an error $\leq$ 0.2s, it should be classified as a TP.
Recall measures the number of correctly predicted positive samples and the FRB search algorithm’s coverage. This measure is critical because it signifies our capacity to identify more FRB signals and indicates the level of completeness of the search. In this paper, we use TP/P to calculate the recall rate.
Precision measures correctly predicted positive samples among all predicted samples. This metric directly affects the judgment of the FRB search’s results, and low precision will result in high labor costs for secondary screening. We use TP/(TP+FP) to calculate the precision rate.
F1 score is a metric for evaluating classification problems. It is the harmonic mean of precision and recall.
$F1=2(\frac{1}{Recall}+\frac{1}{Precision})^{-1}$ (3)
5.2 Results
We run our baseline tasks on a server with 2 Intel(R) Xeon(R) Platinum 8358 CPUs (32 cores 2.6GHz), 1TB memory, and 8 NVIDIA A40 GPUs. PRESTO, Heimdall, and RaSPDAM were deployed to process the dataset with the parameters mentioned above, and the results are presented in Table 2. To process a single sample file of FRB20121102, PRESTO took an average of 141.96 seconds on 1 CPU, Heimdall took an average of 6.98 seconds on 1 GPU, while RaSPDAM only took an average of 3.37 seconds on 1 GPU.
As mentioned in previous sections, different from conventional softwares, RaSPDAM directly converts signal sequences into images rather than spending more computational resources on performing dedispersion. For RaSPDAM, the FITS file only needs to be scanned once, reducing time complexity. Furthermore, in the implementation, the image transformation and model segmentation steps have been accelerated by GPU. Thus, RaSPDAM shows significant speedup in most cases.
Table 2: Methods Running Results: Confusion Matrix and Evaluation Metrics
Software Confusion Matrix Evaluation Metrics
N P N+P TN TP FN FP Recall Precision F1 Score
PRESTO 1000 600 1600 3 472 0 26963700 0.7867 1.7505E-05 3.5009E-05
Heimdall 1000 600 1600 218 489 36 5854 0.8150 0.0771 0.1409
RaSPDAM 1000 600 1600 989 466 128 6 0.7767 0.9873 0.8694
The results indicate that PRESTO and Heimdall have a higher recall rate than the machine learning algorithm RaSPDAM on the total dataset. However, the accuracy of RaSPDAM is very high, whereas both PRESTO and Heimdall are unsatisfactory. Due to the misidentification of RFI as genuine signals, PRESTO and Heimdall calculate a significantly larger number of detected signals than true FRBs. As a result, the absence of human intervention makes it difficult for the two conventional softwares to screen out true FRB signals effectively.
None of the three search pipelines can identify all of the FRB signals according to the final testing results. For conventional software, refining the DM grid and reducing the SNR threshold may improve the recall rate but may lead to even lower precision. For FRB20121102 samples in our dataset, we reduced PRESTO’s DM grid to 0.05 pc $cm^{-3}$ but observed nearly no recall rate changes while the precision rate increases. Nevertheless, raising the SNR threshold in PRESTO to 5.0 decreased the recall rate from 0.7745 to 0.6957 and increased the precision rate from 2.1469E-05 to 0.0110. In conclusion, to tackle the issue observed in PRESTO, it is essential to advance RFI removal methods, which may overlook specific FRB signals during RFI removal. Regarding Heimdall, a more in-depth exploration of the code may be necessary for adjustments due to its limited adjustable parameters. Given that the RaSPDAM algorithm is trained based on simulated signals, a strategy for improvement would be fine-tuning the model using authentic data.
Moreover, as depicted in Figure 7, the performance of the methods varies across different FRB sources due to the inherent differences in FRB attributes among sources. Subsequently, we intend to expand our dataset by incorporating additional sources to enhance its comprehensiveness.
Figure 7: Methods Evaluation Metrics on Different FRB Sources
6 Application
6.1 Distributed Computing Architecture
To address the challenge of processing petabyte-scale data generated by FAST, we adopted a distributed computing architecture based on Kubernetes  [^4] https://kubernetes.io/ and Airflow  [^5] https://airflow.apache.org/. This architecture significantly extends our data processing capabilities, enabling the efficient handling of massive datasets.
  1. 1.
    Data Processing Pipeline: We used Airflow’s Directed Acyclic Graph (DAG) to construct the data processing pipeline. This pipeline includes not only the core algorithm processing but also pre- and post-processing steps such as data cleaning and data organization. It ensures the completeness and continuity of the data processing workflow.
  2. 2.
    Dynamic Task Partitioning: We employed Airflow’s Dynamic Task Mapping feature, which allows for dynamically mapping large data tasks to different processing nodes. Based on the characteristics of the tasks and available resources, tasks are assigned to the most optimal computing nodes. This approach maximizes resource utilization and optimizes task execution efficiency.
6.2 Discoveries of RaSPDAM
The RaSPDAM algorithm has been successfully implemented in FRB and pulsar studies. Since its deployment last year, our team has leveraged this algorithm to identify 2 FRBs (FRB20211103A and FRB20230104) and 80 pulsars. Among these pulsar detections, 13 are previously undiscovered pulsars, highlighting the algorithm’s efficacy in uncovering new celestial objects. Table 3 presents a compilation of these pulsars.
Table 3: Pulsars found by RaSPDAM
No. Pulsar New Discovery No. Pulsar New Discovery
1 19C107_J0528+04 41 PSR J0652+0142
2 19C147_J0104+6127 Yes 42 PSR J0741+17
3 19C148_J0145+6020 Yes 43 PSR J1628+4406
4 19C149_J2114+2630 Yes 44 PSR J1822+2617
5 19C150_J0114+6149 Yes 45 PSR J1829+25
6 19C151_J0000+6250 Yes 46 PSR J1908+2351
7 19C152_J0231+6254 Yes 47 PSR J1911+37
8 19C155_J1958+2332 Yes 48 PSR J1912+2525
9 19C156_J1924+2354 Yes 49 PSR J1919+2621
10 19C158_J1735+3613 Yes 50 PSR J1929+3817
11 19C159_J2131+3642 Yes 51 PSR J1938+2659
12 19C160_J1726+3707 Yes 52 PSR J1939+2449
13 19C171_J0009+5923 Yes 53 PSR J1939+2609
14 19C173_J0642+0238 Yes 54 PSR J1941+2525
15 C10_J0209+2621 55 PSR J1946+2535
16 PSR J0033+61 56 PSR J1946+2611
17 PSR J0058+6125 57 PSR J1946+35
18 PSR J0116+57 58 PSR J1948+2333
19 PSR J0125+62 59 PSR J1948+2551
20 PSR J0139+5814 60 PSR J1954+2529
21 PSR J0141+6009 61 PSR J1956+35
22 PSR J0147+5922 62 PSR J1959+3620
23 PSR J0147+5922 63 PSR J2005+3547
24 PSR J0157+6212 64 PSR J2005+3552
25 PSR J0215+6218 65 PSR J2008+2513
26 PSR J0243+6027 66 PSR J2008+3758
27 PSR J0248+6021 67 PSR J2016+38
28 PSR J0343+06 68 PSR J2019+3718g
29 PSR J0358+4155 69 PSR J2019+3810
30 PSR J0413+58 70 PSR J2021+3651
31 PSR J0435+2749 71 PSR J2026+3656g
32 PSR J0447+04 72 PSR J2027+37
33 PSR J0458+0505 73 PSR J2030+3641
34 PSR J0459+0210 74 PSR J2037+3621
35 PSR J0459+0210 75 PSR J2045+3633
36 PSR J0601+0527 76 PSR J2055+3630
37 PSR J0608+1635 77 PSR J2102+38
38 PSR J0624+0424 78 PSR J2116+3701
39 PSR J0625+17 79 PSR J2156+2618
40 PSR J0627+16
7 Discussion
We provide the FAST-FREX dataset, offer a new machine learning algorithm, RaSPDAM, and use the experiment results of the conventional single-pulse search software PRESTO and Heimdall, along with RaSPDAM as the baseline. The results demonstrate excellent room for progress regarding recall, precision, and performance. Designing and developing new search processes based on machine learning algorithms is a solution. Researchers can also explore machine learning algorithms combined with PRESTO and Heimdall, such as FETCH (Connor & van Leeuwen, 2018), which can reduce the costs of manual screening. The following are some limitations that should be considered when developing machine learning algorithms using the FAST-FREX dataset:
1) a single sample file in our dataset contains either no or one FRB signal. However, most observation files from radio telescopes are significantly longer, containing multiple FRB signals; in some cases, some of the signals may be truncated. When applying newly designed machine learning algorithms to these data, data preprocessing (split or merge) is essential to prevent missing FRB signals at the edges of observation files.
2) Our dataset currently only includes 600 signals from three FRB sources and RFI/noise observed from the specific observations. However, due to variations in observational conditions and intrinsic properties of FRBs, unexpected patterns possibly exist in unknown candidates, which causes models trained on this dataset to fail to recognize other signals. The Blinkverse (Xu et al., 2023) database accumulated an impressive catalog of 8,007 FRB bursts from 813 FRB sources (data up to May 27, 2024), including the bursts our dataset collected. In future work, we aim to extend the dataset with more FRBs, guided by the comprehensive records within Blinkverse, to enhance the comprehensiveness of our dataset.
Furthermore, RaSPDAM can only provide the ToA as a result, while other algorithms provide both ToA and DM, which may cause inconvenience in candidate verification. For future work, we will continue to enhance the capabilities of RaSPDAM to provide more comprehensive features. In addition, in previous model training, we only used simulated FRB signals, which may differ from real signals. We will try to fine-tune the model with real signals in the future, which may further improve the recall rate.
8 Conclusion
This paper presents the FAST-FREX dataset for FRB search based on FAST observation data. Furthermore, we developed a new machine learning algorithm, RaSPDAM, with higher precision, better performance, and a considerable recall rate. Also, we present the benchmark results of conventional FRB search pipelines, PRESTO and Heimdall, which can serve as reference points for future algorithm research.
The dataset aims to support the development of machine learning algorithms for searching FRB signals, such as the method proposed in (Zhang et al., 2024). To ensure comprehensive coverage of the entire parameter space for FRB signals, we plan to collect more FRB signals from observation data acquired by FAST in the future. Moreover, to harness the potential of Artificial Intelligence (AI) for improving the efficiency of pulsar searches and enabling the exploration of pulsars in closed binary systems, we envision the development of a dataset designed for pulsar search tasks.
The dataset is publicly and freely available to the scientific community (https://www.scidb.cn/en/s/VfaYNr). Based on FAST-FREX, a FRB search algorithm challenge will be held to accelerate the research of machine learning algorithms in this field. Meanwhile, we hope that the FAST-FREX dataset serves not only as a facilitator for AI applications in FRB search but also as a catalyst for the advancement of AI for Science.
This work is partially supported by the National Natural Science Foundation of China (NSFC) (11988101). Di Li is a New Cornerstone investigator. This work is also supported by Key R&D Program of Zhejiang (2024SSYS0012). This work made use of the observation data from FAST (Five-hundred-meter Aperture Spherical radio Telescope). FAST is a Chinese national mega-science facility, operated by National Astronomical Observatories, Chinese Academy of Sciences.
Appendix A FAST-FREX Datasheet
In this section, we will comply with the guidelines presented in ”Datasheets for Datasets”  [^6] https://arxiv.org/abs/1803.09010 to provide detailed documentation of the FAST dataset for Fast Radio bursts EXploration (FAST-FREX). Therefore, we will address specific sections that include queries related to the dataset’s motivation, composition, collection process, pre-processing/cleaning/labeling, options for usage, distribution, and maintenance. We will distinguish each section’s primary and subsequent questions using bold and italic fonts.
A.1 Motivation
• For what purpose was the dataset created?
Was there a specific task in mind? Was there a specific gap that needed to be filled?
The FAST-FREX dataset aims to assist researchers in developing advanced machine learning algorithms for searching Fast Radio Burst (FRB) signals. We hope that the FAST-FREX dataset serves not only as a facilitator for Artificial Intelligence (AI) applications in FRB search but also as a catalyst for advancing AI for Science.
• Who created the dataset (e.g., which team, research group) and on behalf of which entity (e.g., company, institution, organization)?
The Research Center for Astronomical Computing of Zhejiang Laboratory created the dataset. The Zhejiang Laboratory and National Astronomical Observatories, Chinese Academy of Sciences (NAOC) are on behalf of the dataset.
• Who funded the creation of the dataset?
If there is an associated grant, please provide the name of the grantor and the grant name and number.
Astronomical Big Data Service Platform Project of the Zhejiang Laboratory funded this dataset. And this work is partially supported by the National Natural Science Foundation of China (NSFC) (11988101). This work is also supported by Key R&D Program of Zhejiang (2024SSYS0012).
• Any other comments?
No.
A.2 Composition
• What do the instances that comprise the dataset represent (e.g., documents, photos, people, countries)?
Are there multiple types of instances (e.g., movies, users, and ratings; people and interactions between them; nodes and edges)?
The FAST-FREX comprises two file types: sample file and parameter file. The sample files, stored in FITS format, contain pre-cropped observation data. The parameter files, stored in CSV format, record various parameters of FRBs in positive samples.
• How many instances are there in total (of each type, if appropriate)?
The dataset contains a total of 1600 instances that are divided into 600 positive sample files containing FRB signals and 1000 negative sample files containing Radio Frequency Interference (RFI) and noise.
• Does the dataset contain all possible instances or is it a sample (not necessarily random) of instances from a larger set?
If the dataset is a sample, then what is the larger set? Is the sample representative of the larger set (e.g., geographic coverage)? If so, please describe how this representativeness was validated/verified. If it is not representative of the larger set, please describe why not (e.g., to cover a more diverse range of instances, because instances were withheld or unavailable).
We carefully chose 600 authentic FRB signals from individual bursts of FRB20121102, FRB20180301, and FRB20201124 as positive examples. To guarantee the diversity of FRB signals in our dataset, we make sure the positive samples have a wide-spread distribution across four dimensions: (1) Full Width at Half Maximum (FWHM), (2) bandwidth, (3) peak flux density ($S_{\rm{peak}}$), and (4) fluence (F). We intend to extend the dataset with more FRB signals to cover the entire parameter space for FRB signals comprehensively.
• What data does each instance consist of?
“Raw” data (e.g., unprocessed text or images) or features? In either case, please provide a description.
Each instance includes observation data in FITS  [^7] https://fits.gsfc.nasa.gov/ format with 60 * 1024 time sampling points, a 98.304 $\mu$m or 49.152 $\mu$s sampling rate, and a single polarization channel. The size of each sample file is approximately 244 MB or 488MB, depending on its sampling rate. Parameter files are stored in CSV format to record various FRB parameters of positive sample files.
• Is there a label or target associated with each instance?
If so, please provide a description.
Yes. Each positive sample file only contains one FRB signal and has a corresponding parameter file. The parameter description file includes the FRB signal’s ToA (Time of Arrival), DM, and MJD. It should be noted that MJD, in this context, refers to surface observation time rather than arrival time at the solar system barycenter after correcting the frequency to 1.5GHz. Negative sample files do not have a corresponding parameter file.
• Is any information missing from individual instances?
If so, please provide a description, explaining why this information is missing (e.g., because it was unavailable). This does not include intentionally removed information, but might include, e.g., redacted text.
We reduced the sample data size by averaging the first two polarization signals while maintaining the signal’s integrity and mitigating the noise’s effect to some extent. The parameter file for the positive samples only includes parameters related to FRB search, such as ToA and DM. Other information, including FWHM, bandwidth, $S_{\rm{peak}}$, and F, are not provided because they are not helpful for FRB search. Instead, these parameters are primarily used for analyzing the detected FRB signals.
• Are relationships between individual instances made explicit (e.g., users’ movie ratings, social network links)?
If so, please describe how these relationships are made explicit.
Yes. The FRB signal in each instance is independent and chosen from the sources FRB20121102, FRB20180301, and FRB20201124.
• Are there recommended data splits (e.g., training, development/validation, testing)?
If so, please provide a description of these splits, explaining the rationale behind them.
No.
• Are there any errors, sources of noise, or redundancies in the dataset?
If so, please provide a description.
Both positive and negative sample files contain noise and RFI because they are inevitable during the Five-hundred-meter Aperture Spherical radio Telescope (FAST) observations.
• Is the dataset self-contained, or does it link to or otherwise rely on external resources (e.g., websites, tweets, other datasets)?
If it links to or relies on external resources, a) are there guarantees that they will exist, and remain constant, over time; b) are there official archival versions of the complete dataset (i.e., including the external resources as they existed at the time the dataset was created); c) are there any restrictions (e.g., licenses, fees) associated with any of the external resources that might apply to a dataset consumer? Please provide descriptions of all external resources and any restrictions associated with them, as well as links or other access points, as appropriate.
Our dataset is entirely self-contained and comprises samples collected from publicly available FAST observation data  [^8] https://fast.bao.ac.cn/.
• Does the dataset contain data that might be considered confidential (e.g., data that is protected by legal privilege or by doctor-patient confidentiality, data that includes the content of individuals’ non-public communications)?
If so, please provide a description.
No.
• Does the dataset contain data that, if viewed directly, might be offensive, insulting, threatening, or might otherwise cause anxiety?
If so, please describe why.
No.
A.3 Collection Process
• How was the data associated with each instance acquired?
Was the data directly observable (e.g., raw text, movie ratings), reported by subjects (e.g., survey responses), or indirectly inferred/derived from other data (e.g., part-of-speech tags, model-based guesses for age or language)? If the data was reported by subjects or indirectly inferred/derived from other data, was the data validated/verified? If so, please describe how.
We build the dataset using observational data from reports on FRB20121102  [^9] https://www.nature.com/articles/s41586-021-03878-5, FRB20180301  [^10] https://iopscience.iop.org/article/10.3847/1538-4357/ac63a8 and FRB20201124  [^11] http://groups.bao.ac.cn/ism/CRAFTS/FRB20201124A/, which contains detailed information about individual bursts observed by FAST.
• What mechanisms or procedures were used to collect the data (e.g., hardware apparatuses or sensors, manual human curation, software programs, software APIs)?
How were these mechanisms or procedures validated?
The raw observation data were collected from three FRB sources. The FAST continuous monitoring of FRB121102 began in August 2019. During 59.5 hours, 1,652 independent burst events were detected from August 29th to October 29th. Meanwhile, FAST observed FRB20180301 on 2021 March 4th, 9th, and 19th and detected five bright radio bursts. 881 bursts detected from FRB20201124 by FAST between 2021 September 25th and 28th. The data was gathered using 4096 frequency channels over 1.05 GHz to 1.45 GHz, with 0.122 MHz frequency resolution. FRB20121102 has a 98.304 $\mu$s sampling rate, while others have a 49.152 $\mu$s sampling rate. These channels recorded four polarization signals.
To preserve the integrity of FRB signals, we set the cropping duration of each positive sample file to about 6 seconds. We reduced the data size by averaging the first two polarization signals while maintaining complete signal recordings and mitigating the effect of noise to some extent. We carefully chose 600 bursts to create positive samples. After eliminating the observation data of FRB signals, we extracted 1000 negative samples of RFI and noise with the same data length as the positive samples.
• If the dataset is a sample from a larger set, what was the sampling strategy (e.g., deterministic, probabilistic with specific sampling probabilities)?
This dataset is a subset of bursts from three FRB sources. Positive samples have been manually chosen and guaranteed to exhibit changes in FWHM, bandwidth, $S_{\rm{peak}}$, and F four dimensions. Figure 1 in the paper depicts the probability distribution of these four parameters.
• Who was involved in the data collection process (e.g., students, crowdworkers, contractors) and how were they compensated (e.g., how much were crowdworkers paid)?
The official staff of the Zhejiang Laboratory and NAOC collected the dataset.
• Over what timeframe was the data collected?
Does this timeframe match the creation timeframe of the data associated with the instances (e.g., recent crawl of old news articles)? If not, please describe the time- frame in which the data associated with the instances was created.
The raw observation data for FRB121102 was generated from August 29th to October 29th, 2019, within 59.5 hours. FRB20180301 was observed on 2021 March 4th, 9th, and 19th, and FRB20201124 was between 2021 September 25th and 28th. However, the dataset creation process was completed within a relatively short period.
• Were any ethical review processes conducted (e.g., by an institutional review board)?
If so, please provide a description of these review processes, including the outcomes, as well as a link or other access point to any supporting documentation.
No.
A.4 Preprocessing/Cleaning/Labeling
• Was any preprocessing/cleaning/labeling of the data done (e.g., discretization or bucketing, tokenization, part-of-speech tagging, SIFT feature extraction, removal of instances, processing of missing values)?
If so, please provide a description. If not, you may skip the remaining questions in this section.
To maintain the integrity of FRB signals, the cropping duration of each positive sample file was set to 6 seconds. We also ensure that the events appear randomly within the observation time covered by the file rather than setting a fixed ToA. This approach simulates FRB detection more realistically and enhances the dataset’s diversity, which is beneficial for algorithm research. We reduced the data size by averaging the first two polarization signals while maintaining the signal’s integrity and mitigating the noise’s effect to some extent. Each sample file consists of 60 * 1024 or 120 * 1024 time sampling points, and the number of polarization channels is reduced to one, which differs from the original data. The size of each file is approximately 244 MB or 488MB, depending on its sampling rate.
• Was the “raw” data saved in addition to the preprocessed/cleaned/labeled data (e.g., to support unanticipated future uses)?
If so, please provide a link or other access point to the “raw” data.
No. Due to the large amount of raw data, it is not convenient to store.
• Is the software that was used to preprocess/clean/label the data available?
If so, please provide a link or other access point.
No, but we have explained the dataset creation approach in the preceding sections.
• Any other comments?
No.
A.5 Uses
• Has the dataset been used for any tasks already?
If so, please provide a description.
We have performed our experiments using conventional single-pulse search softwares, PRESTO and Heimdall, and a new machine learning algorithm Radio Single-Pulse Detection Algorithm Based on Visual Morphological Features (RaSPDAM), as outlined in Section 5 of this paper.
• Is there a repository that links to any or all papers or systems that use the dataset?
If so, please provide a link or other access point.
No.
• What (other) tasks could the dataset be used for?
The dataset can be used to train and test machine learning algorithms for FRB search.
• Is there anything about the composition of the dataset or the way it was collected and preprocessed/cleaned/labeled that might impact future uses?
For example, is there anything that a dataset consumer might need to know to avoid uses that could result in unfair treatment of individuals or groups (e.g., stereotyping, quality of service issues) or other risks or harms (e.g., legal risks, financial harms)? If so, please provide a description. Is there anything a dataset consumer could do to mitigate these risks or harms?
No.
• Are there tasks for which the dataset should not be used?
If so, please provide a description.
No.
• Any other comments?
No.
A.6 Distribution
• Will the dataset be distributed to third parties outside of the entity (e.g., company, institution, organization) on behalf of which the dataset was created?)
If so, please provide a description.
Yes, the dataset is openly distributed to third parties outside the entity on behalf of which the dataset was created. It is an open-source dataset for anyone to access, use, and share for noncommercial purposes.
• How will the dataset will be distributed (e.g., tarball on website, API, GitHub)?
Does the dataset have a digital object identifier (DOI)?
The URL to access the dataset is https://www.scidb.cn/en/s/VfaYNr.
• When will the dataset be distributed?
The dataset is publicly accessible.
• Will the dataset be distributed under a copyright or other intellectual property (IP) license, and/or under applicable terms of use (ToU)?
If so, please describe this license and/or ToU, and provide a link or other access point to, or otherwise reproduce, any relevant licensing terms or ToU, as well as any fees associated with these restrictions.
The dataset is openly shared under the CC BY-NC-ND 4.0 license.
• Have any third parties imposed IP-based or other restrictions on the data associated with the instances?
If so, please describe these restrictions, and provide a link or other access point to, or otherwise reproduce, any relevant licensing terms, as well as any fees associated with these restrictions.
No.
• Do any export controls or other regulatory restrictions apply to the dataset or to individual instances?
If so, please describe these restrictions, and provide a link or other access point to, or otherwise reproduce, any supporting documentation.
No.
• Any other comments?
No
A.7 Maintenance
• Who will be supporting/hosting/maintaining the dataset?
The Zhejiang Laboratory will host and maintain the dataset.
• How can the owner/curator/manager of the dataset be contacted (e.g., email address)?
Please reach out to Huaxi Chen (chenhuaxi@zhejianglab.com) and Xuerong Guo (guoxr@zhejianglab.com), the authors of this paper.
• Is there an erratum?
If so, please provide a link or other access point.
No.
• Will the dataset be updated (e.g., to correct labeling errors, add new instances, delete instances)?
If so, please describe how often, by whom, and how updates will be communicated to dataset consumers (e.g., mailing list, GitHub)?
Our dataset will be periodically updated to add more samples from various FRB sources. Subsequent modifications will be reported on its public website.
• If the dataset relates to people, are there applicable limits on the retention of the data associated with the instances (e.g., were the individuals in question told that their data would be retained for a fixed period of time and then deleted)?
If so, please describe these limits and explain how they will be enforced.
No.
• Will older versions of the dataset continue to be supported/hosted/maintained?
If so, please describe how. If not, please describe how its obsolescence will be communicated to dataset consumers.
Yes, information on updates and maintenance will be regularly posted on its public website.
• If others want to extend/augment/build on/contribute to the dataset, is there a mechanism for them to do so?
If so, please provide a description. Will these contributions be validated/verified? If so, please describe how. If not, why not? Is there a process for communicating/distributing these contributions to dataset consumers? If so, please provide a description.
No.
• Any other comments?
No
A.8 Responsibility
The authors bear all responsibility for rights violations of the FAST-FREX dataset.
Appendix B Reading and Using the Dataset
The sample files are stored in the standard FITS format and can be accessed through the Python library astropy. These files comprise two Header and Data Unit (HDU) components. Observation parameters are stored in the headers of the first and second HDU, while observation data are stored in the second HDU. The observation data contains five dimensions: 1) the first two dimensions consist of 60 * 1024 or 120* 1024 time sampling points, 2) the third dimension represents one polarization, 3) the fourth dimension includes 4096 frequency channels, and 4) the fifth dimension is currently unused.
Please note that because the sample files were generated from the raw observation files, the start MJD time parameters do not apply to the current sample files, particularly the STT_IMJD, STT_SMJD, and STT_OFFS parameters in the first header (f[0].header in the following code).
Please refer to the code example below for specifics.
1from astropy.io import fits
2
3with fits.open(’./FRB20121102/FRB20121102_0001.fits’) as f:
4 # File structure information.
5 f.info()
6
7 # f[1].header: observation parameters
[…]
9 print("Time sampling points.: {} * {}".format(f[1].header[’NAXIS2’], f[1].header[’NSBLK’]))
10 print("Nr of polarisations: {}".format(f[1].header[’NPOL’]))
11 print("Number of channels/sub-bands in this file: {}".format(f[1].header[’NCHAN’]))
12
13 # f[1].data: observation data
[…]
15
16 # output:
17 # No. Name Ver Type Cards Dimensions Format
18 # 0 PRIMARY 1 PrimaryHDU 59 ()
19 # 1 SUBINT 1 BinTableHDU 76 60R x 17C [1D, 1D, 1D, 1D, 1D, 1D, 1D, 1E, 1E, 1E, 1E, 1E, 4096E, 4096E, 4096E, 4096E, 4194304B]
20 #
21 # Time per bin or sample: 9.8304e-05
22 # Time sampling points.: 60 * 1024
23 # Nr of polarisations: 1
24 # Number of channels/sub-bands in this file: 4096
25 #
26 # data shape: (60, 1024, 1, 4096, 1)
Appendix C Baseline Models Experiments
The paper has provided a detailed explanation of the pipelines of these two conventional single-pulse search softwares, PRESTO and Heimdall, and machine learning algorithm RaSPDAM. Here, we will provide the commands and parameters executed in our experiments.
C.1 PRESTO
The software repository link for conventional single-pulse search software PRESTO is https://github.com/scottransom/presto.
1# RFI Removal
2rfifind FRB20121102_0001.fits -o test1 -time 1
3
4# De-dispersion
5prepsubband -nobary -numout 524288 -nsub 1024 -lodm 350.0 -dmstep 1 -numdms 300 -downsamp 1 -mask test1_rfifind.mask -o FRB20121102_0001 FRB20121102_0001.fits
6
7# Single-pulse Search
8ls *.dat |xargs -n 408 -P 12 python single_pulse_search.py -b -m 2 -t 3.0
C.2 Heimdall
The software repository link for conventional single-pulse search software Heimdall is https://sourceforge.net/projects/heimdall-astro/. We used your_heimdall  [^12] https://github.com/thepetabyteproject/your/blob/main/bin/your_heimdall.py to test Heimdall on our dataset, as Heimdall’s source code does not support processing FITS files and requires a file format conversion. Your_heimdall can complete such conversion, essentially using Heimdall to process the file.
1
2your_heimdall.py -dm 350 650 -g 0 -f FRB20121102_0001.fits
C.3 RaSPDAM
The RaSPDAM was implemented in Python. The source code and model file can be found in the attachment. The following example demonstrates an evaluation using a FITS file.
1
2python3 slope_detection.py -o results FRB20121102_0001.fits
Appendix D Dataset Positive Samples Parameters
Table 4: Parameters of 600 Positive Samples in Dataset
Burst File Name Source MJD DM ToA FWHM Bandwidth Peak Flux Density Fluence
(pc $cm^{-3}$) (s) (ms) (MHz) (Jy) (Jy ms)
FRB20121102_0001.fits FRB20121102 58724.87756 567.3 3.0926 6.13 110 0.012831 0.0787
FRB20121102_0002.fits FRB20121102 58724.88513 565.1 3.3074 2.83 400 0.26905 0.7614
FRB20121102_0003.fits FRB20121102 58724.88637 565.8 1.0731 8.51 160 0.00627 0.0534
FRB20121102_0004.fits FRB20121102 58724.88758 567.6 4.1216 5.49 200 0.01248 0.0685
FRB20121102_0005.fits FRB20121102 58724.8911 565.8 1.2662 7.99 230 0.00794 0.0635
FRB20121102_0006.fits FRB20121102 58724.93325 564.1 4.1918 2.01 320 0.0585 0.1176
FRB20121102_0007.fits FRB20121102 58724.94249 568.1 3.0237 5.41 230 0.01315 0.0711
FRB20121102_0008.fits FRB20121102 58724.95614 568 3.5371 4.6 190 0.01413 0.065
FRB20121102_0009.fits FRB20121102 58724.97639 564 4.4543 5.19 190 0.01614 0.0838
FRB20121102_0010.fits FRB20121102 58725.87599 567.1 1.2409 2.85 250 0.13023 0.3712
FRB20121102_0011.fits FRB20121102 58725.87621 562.3 0.6020 6.13 210 0.00446 0.0274
FRB20121102_0012.fits FRB20121102 58725.88178 565.9 0.2708 6.26 290 0.0065 0.0407
FRB20121102_0013.fits FRB20121102 58725.88564 566 2.0624 3.84 150 0.01454 0.0559
FRB20121102_0014.fits FRB20121102 58725.88943 564.5 2.0019 2.03 100 0.14407 0.2925
FRB20121102_0015.fits FRB20121102 58725.89072 566.1 3.2988 4.44 100 0.06265 0.2782
FRB20121102_0016.fits FRB20121102 58725.89318 564.5 4.3049 3.12 100 0.09759 0.3045
FRB20121102_0017.fits FRB20121102 58725.89696 559.7 0.4314 3.66 180 0.02089 0.0766
FRB20121102_0018.fits FRB20121102 58725.89877 568.5 3.3153 8.88 130 0.01146 0.1017
FRB20121102_0019.fits FRB20121102 58725.90072 566.7 4.4214 4.92 130 0.01306 0.0642
FRB20121102_0020.fits FRB20121102 58725.91554 568.5 2.9621 5.89 200 0.01822 0.1073
FRB20121102_0021.fits FRB20121102 58725.91635 566 2.3250 3.38 120 0.0094 0.0318
FRB20121102_0022.fits FRB20121102 58725.91764 566 3.5363 3.37 105 0.01415 0.0478
FRB20121102_0023.fits FRB20121102 58725.9189 566 1.7267 3.54 200 0.01022 0.0362
FRB20121102_0024.fits FRB20121102 58725.92026 566.3 3.3744 1.92 150 0.15473 0.2971
FRB20121102_0025.fits FRB20121102 58725.92985 563.2 2.0739 4.54 220 0.00919 0.0417
FRB20121102_0026.fits FRB20121102 58725.93131 567.6 1.3954 7.91 290 0.01508 0.1192
FRB20121102_0027.fits FRB20121102 58725.93237 566 4.0378 3.81 300 0.01373 0.0523
FRB20121102_0028.fits FRB20121102 58725.93294 569.4 0.8811 2.72 140 0.0085 0.0231
FRB20121102_0029.fits FRB20121102 58725.94049 569.4 0.5990 3.52 280 0.01428 0.0503
FRB20121102_0030.fits FRB20121102 58725.9409 566 3.9965 4.04 180 0.0125 0.0505
FRB20121102_0031.fits FRB20121102 58725.94223 566 3.0291 2.9 210 0.02341 0.068
FRB20121102_0032.fits FRB20121102 58725.95399 565.7 4.5690 5.45 200 0.07952 0.4334
FRB20121102_0033.fits FRB20121102 58725.95541 554.5 1.0340 4.31 160 0.00709 0.0306
FRB20121102_0034.fits FRB20121102 58725.9566 567 0.0760 3.47 100 0.06263 0.2173
FRB20121102_0035.fits FRB20121102 58725.9583 569 1.1831 5.89 150 0.05868 0.3456
FRB20121102_0036.fits FRB20121102 58725.96176 570 1.6019 10.3 100 0.03169 0.3264
FRB20121102_0037.fits FRB20121102 58725.96432 565.9 4.2382 9.14 200 0.10011 0.915
FRB20121102_0038.fits FRB20121102 58725.9665 564.6 2.3964 2.55 300 0.24736 0.6308
FRB20121102_0039.fits FRB20121102 58725.97352 566 0.3103 4.06 200 0.01802 0.0732
FRB20121102_0040.fits FRB20121102 58725.97785 566 1.6335 4.03 200 0.01309 0.0527
FRB20121102_0041.fits FRB20121102 58725.98146 567.9 2.8102 8.52 250 0.03788 0.3227
FRB20121102_0042.fits FRB20121102 58725.98847 566 2.1579 3.75 100 0.00924 0.0346
FRB20121102_0043.fits FRB20121102 58725.99225 563.7 0.8485 3.19 140 0.00901 0.0288
FRB20121102_0044.fits FRB20121102 58725.99259 566 2.7917 2.67 150 0.01783 0.0475
FRB20121102_0045.fits FRB20121102 58725.99453 563 2.1191 5.33 50 0.04107 0.2189
FRB20121102_0046.fits FRB20121102 58725.9983 570 0.8389 5.52 100 0.0475 0.2622
FRB20121102_0047.fits FRB20121102 58726.9185 564.6 3.3889 2.6 200 0.14299 0.3718
FRB20121102_0048.fits FRB20121102 58726.92158 568.9 4.1755 7.53 170 0.0083 0.0625
FRB20121102_0049.fits FRB20121102 58726.92408 566.3 0.6134 4.45 330 0.0126 0.0561
FRB20121102_0050.fits FRB20121102 58726.92772 550.6 2.6145 4.44 65 0.01136 0.0505
FRB20121102_0051.fits FRB20121102 58726.92802 564.5 3.4432 6.96 120 0.01402 0.0975
FRB20121102_0052.fits FRB20121102 58726.93712 566 1.4909 2.88 170 0.00732 0.0211
FRB20121102_0053.fits FRB20121102 58726.94971 565.6 0.7280 4.64 50 0.04632 0.2149
FRB20121102_0054.fits FRB20121102 58726.96581 566.2 0.5840 3.01 100 0.0511 0.1538
FRB20121102_0055.fits FRB20121102 58726.96834 567.9 0.0790 5.23 300 0.07034 0.3679
FRB20121102_0056.fits FRB20121102 58726.99613 568.5 4.2030 3.74 195 0.00414 0.0155
FRB20121102_0057.fits FRB20121102 58727.00019 566 3.0518 3.85 180 0.00893 0.0343
FRB20121102_0058.fits FRB20121102 58727.00142 566 3.3179 4.11 100 0.01053 0.0433
FRB20121102_0059.fits FRB20121102 58727.00404 566 4.4402 3.99 180 0.01041 0.0415
FRB20121102_0060.fits FRB20121102 58727.00522 569.4 0.9801 3.35 75 0.00256 0.0086
FRB20121102_0061.fits FRB20121102 58727.01219 568.5 0.2050 7.54 160 0.01059 0.0799
FRB20121102_0062.fits FRB20121102 58727.01422 566 1.1606 3.72 110 0.01279 0.0476
FRB20121102_0063.fits FRB20121102 58727.02681 565.7 0.0360 4.75 350 0.10944 0.5198
FRB20121102_0064.fits FRB20121102 58727.02699 566 0.6551 3.98 250 0.01353 0.0538
FRB20121102_0065.fits FRB20121102 58727.0296 566 2.8529 3.58 170 0.01324 0.0474
FRB20121102_0066.fits FRB20121102 58727.04422 566 2.9126 2.1 150 0.01969 0.0413
FRB20121102_0067.fits FRB20121102 58727.07798 563.2 1.9210 5.23 30 0.0087 0.0455
FRB20121102_0068.fits FRB20121102 58727.08786 564.5 0.8655 6.56 150 0.01417 0.0929
FRB20121102_0069.fits FRB20121102 58727.09232 565.8 1.9720 10.94 400 0.04737 0.5182
FRB20121102_0070.fits FRB20121102 58727.09706 567.6 0.5870 8.82 165 0.00875 0.0772
FRB20121102_0071.fits FRB20121102 58727.90379 566 1.8139 4.04 460 0.09887 0.3996
FRB20121102_0072.fits FRB20121102 58727.90656 565 1.7979 4.75 140 0.00606 0.0288
FRB20121102_0073.fits FRB20121102 58727.91153 567.2 0.3777 2.62 110 0.00483 0.0127
FRB20121102_0074.fits FRB20121102 58727.91526 567.2 0.2896 8.12 90 0.0048 0.039
FRB20121102_0075.fits FRB20121102 58727.91574 563.8 2.8607 2.81 300 0.5953 1.6728
FRB20121102_0076.fits FRB20121102 58727.92178 566 2.8659 3.83 190 0.03688 0.1411
FRB20121102_0077.fits FRB20121102 58727.92596 561 3.3395 4 170 0.00746 0.0298
FRB20121102_0078.fits FRB20121102 58727.92854 562.8 1.3202 3.78 100 0.09252 0.3497
FRB20121102_0079.fits FRB20121102 58727.93819 566 2.6919 4.55 175 0.01228 0.0559
FRB20121102_0080.fits FRB20121102 58727.95377 564.5 0.2191 3.15 102 0.00722 0.0227
FRB20121102_0081.fits FRB20121102 58727.9627 569 1.1860 4.03 105 0.01444 0.0582
FRB20121102_0082.fits FRB20121102 58727.96275 567.6 0.1072 5.46 110 0.00989 0.054
FRB20121102_0083.fits FRB20121102 58727.96395 566 1.6053 3.13 210 0.01259 0.0394
FRB20121102_0084.fits FRB20121102 58727.96483 552.4 2.6858 7.35 175 0.00482 0.0354
FRB20121102_0085.fits FRB20121102 58727.96512 566 1.4534 4.37 130 0.01041 0.0455
FRB20121102_0086.fits FRB20121102 58727.97165 565.5 0.9240 3.86 200 0.06539 0.2524
FRB20121102_0087.fits FRB20121102 58727.98961 565.5 2.5525 6.47 350 0.70504 4.5616
FRB20121102_0088.fits FRB20121102 58727.99693 566.7 1.2139 6.47 130 0.00498 0.0322
FRB20121102_0089.fits FRB20121102 58727.99986 565.6 3.6556 3.5 50 0.04882 0.1709
FRB20121102_0090.fits FRB20121102 58728.00245 566.1 2.2279 3.38 200 0.04817 0.1628
FRB20121102_0091.fits FRB20121102 58728.00686 562.6 2.3000 5.93 400 0.10736 0.6366
FRB20121102_0092.fits FRB20121102 58728.01267 566.6 2.1056 8.9 100 0.03219 0.2865
FRB20121102_0093.fits FRB20121102 58728.01462 569 0.1214 5.42 125 0.00457 0.0248
FRB20121102_0094.fits FRB20121102 58728.02096 565.1 3.4972 3.71 200 0.7118 2.6408
FRB20121102_0095.fits FRB20121102 58728.02325 563.7 0.4322 6.2 180 0.00949 0.0588
FRB20121102_0096.fits FRB20121102 58728.0269 558.9 0.4844 8.52 400 0.00169 0.0144
FRB20121102_0097.fits FRB20121102 58728.05237 566 3.2574 2.9 300 0.0127 0.0369
FRB20121102_0098.fits FRB20121102 58728.0573 570 2.5815 9.26 250 0.10363 0.9596
FRB20121102_0099.fits FRB20121102 58728.0691 566 4.1637 3.03 400 0.01095 0.0332
FRB20121102_0100.fits FRB20121102 58728.07772 567 3.4837 3.53 200 0.09654 0.3408
FRB20121102_0101.fits FRB20121102 58728.98495 564.3 2.2867 2.69 50 0.08237 0.2216
FRB20121102_0102.fits FRB20121102 58728.98842 567.2 3.2979 6.04 150 0.08738 0.5278
FRB20121102_0103.fits FRB20121102 58728.99392 566 4.6156 3.24 200 0.18703 0.606
FRB20121102_0104.fits FRB20121102 58729.00494 568.3 4.4047 4.41 400 0.0647 0.2853
FRB20121102_0105.fits FRB20121102 58729.0085 569 0.5212 6.05 50 0.03638 0.2201
FRB20121102_0106.fits FRB20121102 58729.01463 568 2.3417 4.05 150 0.05107 0.2068
FRB20121102_0107.fits FRB20121102 58729.01705 564.6 0.2937 2.02 50 0.0867 0.1751
FRB20121102_0108.fits FRB20121102 58729.01805 564.7 1.9315 2.88 400 0.0101 0.0291
FRB20121102_0109.fits FRB20121102 58729.01846 563.1 4.4809 2.2 200 0.16156 0.3554
FRB20121102_0110.fits FRB20121102 58729.02395 565.4 1.2113 3.15 100 0.07467 0.2352
FRB20121102_0111.fits FRB20121102 58729.02706 565.1 0.1473 5.27 150 0.04704 0.2479
FRB20121102_0112.fits FRB20121102 58729.04583 566.6 0.5030 6.79 150 0.05874 0.3988
FRB20121102_0113.fits FRB20121102 58729.04737 563.5 2.5113 2.21 200 0.11853 0.262
FRB20121102_0114.fits FRB20121102 58729.05123 563.8 3.7266 3.44 300 0.1523 0.5239
FRB20121102_0115.fits FRB20121102 58729.05518 568.4 1.6226 5.63 200 0.04226 0.2379
FRB20121102_0116.fits FRB20121102 58729.05609 561.5 3.0189 5.62 190 0.00628 0.0352
FRB20121102_0117.fits FRB20121102 58729.06361 569.4 2.9349 8.41 150 0.06767 0.5691
FRB20121102_0118.fits FRB20121102 58729.06916 569.5 3.0995 4.12 150 0.18388 0.7576
FRB20121102_0119.fits FRB20121102 58729.07307 564.9 3.0734 2.1 300 0.58318 1.2247
FRB20121102_0120.fits FRB20121102 58730.87132 568.1 3.4382 7.98 100 0.00909 0.0725
FRB20121102_0121.fits FRB20121102 58730.87347 567 1.2238 6.88 150 0.05711 0.3929
FRB20121102_0122.fits FRB20121102 58730.88025 569.5 2.2721 8.43 50 0.0387 0.3262
FRB20121102_0123.fits FRB20121102 58730.88097 568.1 1.6585 1.79 200 0.00723 0.0129
FRB20121102_0124.fits FRB20121102 58730.89092 568.9 1.0673 10.11 200 0.04934 0.4988
FRB20121102_0125.fits FRB20121102 58730.89453 566.9 0.5608 3.56 250 0.10085 0.359
FRB20121102_0126.fits FRB20121102 58730.90048 566.2 0.6540 4.05 250 0.10584 0.4287
FRB20121102_0127.fits FRB20121102 58730.90213 567 1.5877 4.89 400 0.07947 0.3886
FRB20121102_0128.fits FRB20121102 58730.90753 563.6 4.1419 2.71 200 0.44693 1.2112
FRB20121102_0129.fits FRB20121102 58730.90783 566.9 1.9505 7.29 200 0.04506 0.3285
FRB20121102_0130.fits FRB20121102 58732.86875 562.8 1.1306 5.94 90 0.00696 0.0414
FRB20121102_0131.fits FRB20121102 58732.8689 566 2.2367 3.76 170 0.01011 0.0381
FRB20121102_0132.fits FRB20121102 58732.87047 566 2.1816 3.54 170 0.01775 0.0627
FRB20121102_0133.fits FRB20121102 58732.87607 566 0.7087 3.13 105 0.0186 0.0582
FRB20121102_0134.fits FRB20121102 58732.87836 563.7 2.3886 7.2 80 0.00608 0.0438
FRB20121102_0135.fits FRB20121102 58732.87877 569.4 1.6878 8.56 140 0.00898 0.0768
FRB20121102_0136.fits FRB20121102 58732.95231 565.9 1.0176 7.53 130 0.00778 0.0586
FRB20121102_0137.fits FRB20121102 58732.95423 566.7 0.0672 2.96 110 0.00692 0.0205
FRB20121102_0138.fits FRB20121102 58732.95433 564.1 0.6848 5.71 200 0.00954 0.0544
FRB20121102_0139.fits FRB20121102 58732.95621 566 2.9801 2.36 130 0.01475 0.0348
FRB20121102_0140.fits FRB20121102 58732.95934 567.2 2.5572 7.3 160 0.00721 0.0526
FRB20121102_0141.fits FRB20121102 58732.96003 568.4 3.8131 4.81 100 0.08157 0.3924
FRB20121102_0142.fits FRB20121102 58732.96073 565.4 0.5333 3.53 120 0.01096 0.0387
FRB20121102_0143.fits FRB20121102 58732.96268 565.4 0.0903 4.14 210 0.39424 1.6337
FRB20121102_0144.fits FRB20121102 58732.96286 566 3.9247 2.83 130 0.01545 0.0437
FRB20121102_0145.fits FRB20121102 58732.96467 567.6 3.1886 4.18 100 0.00914 0.0382
FRB20121102_0146.fits FRB20121102 58732.96766 566 3.0219 4.61 160 0.01191 0.0549
FRB20121102_0147.fits FRB20121102 58732.9727 566 4.0211 3.17 30 0.01006 0.0319
FRB20121102_0148.fits FRB20121102 58732.9731 566 0.3464 1.85 120 0.01839 0.034
FRB20121102_0149.fits FRB20121102 58732.97436 566 1.0993 2.18 140 0.01223 0.0266
FRB20121102_0150.fits FRB20121102 58732.97558 566 1.2829 4.04 130 0.01217 0.0492
FRB20121102_0151.fits FRB20121102 58732.97641 565 2.6086 6.6 130 0.00957 0.0632
FRB20121102_0152.fits FRB20121102 58732.97759 567.2 0.8794 78.52 170 0.00999 0.7844
FRB20121102_0153.fits FRB20121102 58732.97994 567.3 4.2107 6.15 250 0.0741 0.4557
FRB20121102_0154.fits FRB20121102 58732.98207 566 1.4323 3.73 260 0.01173 0.0438
FRB20121102_0155.fits FRB20121102 58732.98576 567.2 4.1596 3.29 100 0.06227 0.2049
FRB20121102_0156.fits FRB20121102 58732.98655 559.7 0.9680 2.4 70 0.00519 0.0125
FRB20121102_0157.fits FRB20121102 58732.99195 565.4 0.6422 7.54 90 0.00767 0.0579
FRB20121102_0158.fits FRB20121102 58733.90523 564.5 0.4174 2.67 400 0.84239 2.2492
FRB20121102_0159.fits FRB20121102 58733.91736 565.9 4.3268 0.34 105 0.00474 0.0016
FRB20121102_0160.fits FRB20121102 58733.91851 566.1 3.6670 2.65 100 0.07565 0.2005
FRB20121102_0161.fits FRB20121102 58733.91876 567.9 0.0511 5.55 250 0.06527 0.3623
FRB20121102_0162.fits FRB20121102 58733.91948 565.5 3.3249 2.26 120 0.09033 0.2042
FRB20121102_0163.fits FRB20121102 58733.92254 562.6 2.9314 1.08 50 0.0421 0.0455
FRB20121102_0164.fits FRB20121102 58733.92316 564.2 0.4492 1.83 100 0.07733 0.1415
FRB20121102_0165.fits FRB20121102 58733.93106 566 3.2304 3.94 180 0.00841 0.0331
FRB20121102_0166.fits FRB20121102 58733.93123 565.6 0.0597 4.61 70 0.01811 0.0836
FRB20121102_0167.fits FRB20121102 58735.01934 564.4 3.1493 5.92 200 0.07251 0.4293
FRB20121102_0168.fits FRB20121102 58735.02504 568.4 4.4950 3.66 230 0.05453 0.1996
FRB20121102_0169.fits FRB20121102 58735.02663 566.7 0.9791 21.27 400 0.0311 0.6615
FRB20121102_0170.fits FRB20121102 58735.02707 563.9 0.4344 2.58 350 0.67999 1.7544
FRB20121102_0171.fits FRB20121102 58735.02846 566.1 0.6257 5.38 180 0.08944 0.4812
FRB20121102_0172.fits FRB20121102 58735.02903 562.8 2.6206 6.74 130 0.00745 0.0502
FRB20121102_0173.fits FRB20121102 58735.03188 562.6 3.5275 1.99 100 0.09554 0.1901
FRB20121102_0174.fits FRB20121102 58735.03225 567.6 3.0726 4.07 150 0.09726 0.3959
FRB20121102_0175.fits FRB20121102 58735.03723 567.4 2.1610 5.02 200 0.07755 0.3893
FRB20121102_0176.fits FRB20121102 58735.04017 566.5 4.0428 4.49 300 0.06999 0.3143
FRB20121102_0177.fits FRB20121102 58735.04043 568.3 2.0596 11.01 300 0.04435 0.4883
FRB20121102_0178.fits FRB20121102 58735.04416 563 4.6303 3.71 250 0.09052 0.3358
FRB20121102_0179.fits FRB20121102 58735.04608 565 3.2099 7.5 132 0.00654 0.0491
FRB20121102_0180.fits FRB20121102 58736.96951 569.5 2.4488 8.85 110 0.04439 0.3929
FRB20121102_0181.fits FRB20121102 58736.97536 566.5 0.9557 4.58 450 0.04668 0.2138
FRB20121102_0182.fits FRB20121102 58736.98094 564.3 4.4411 2.86 300 0.32433 0.9276
FRB20121102_0183.fits FRB20121102 58736.98189 567.2 1.6646 7.33 400 0.04762 0.3491
FRB20121102_0184.fits FRB20121102 58736.98387 566.7 0.6317 4.19 120 0.05879 0.2463
FRB20121102_0185.fits FRB20121102 58736.98432 566.7 3.0716 7.06 150 0.05033 0.3553
FRB20121102_0186.fits FRB20121102 58736.98563 567.7 1.6761 5.03 250 0.04024 0.2024
FRB20121102_0187.fits FRB20121102 58736.98838 569.8 0.9442 6.87 100 0.04001 0.2749
FRB20121102_0188.fits FRB20121102 58736.99206 563.6 3.8486 3.52 350 0.13712 0.4827
FRB20121102_0189.fits FRB20121102 58736.99293 563.5 0.3799 2.99 150 0.08017 0.2397
FRB20121102_0190.fits FRB20121102 58738.95952 566 3.9626 4.49 200 0.01015 0.0456
FRB20121102_0191.fits FRB20121102 58738.96014 559.7 1.9365 6.61 170 0.0104 0.0688
FRB20121102_0192.fits FRB20121102 58738.96335 565.8 0.2539 3.95 50 0.00681 0.0269
FRB20121102_0193.fits FRB20121102 58738.96367 565.8 1.1001 5.77 50 0.00734 0.0424
FRB20121102_0194.fits FRB20121102 58738.96742 567.1 1.5596 8.61 100 0.00453 0.039
FRB20121102_0195.fits FRB20121102 58738.96876 565.1 1.0335 6.79 50 0.00912 0.0619
FRB20121102_0196.fits FRB20121102 58738.97076 568.5 4.2439 3.45 50 0.0077 0.0266
FRB20121102_0197.fits FRB20121102 58738.97434 564.8 4.4162 4.06 200 0.00792 0.0322
FRB20121102_0198.fits FRB20121102 58738.97624 564.9 2.5251 1.91 50 0.01044 0.0199
FRB20121102_0199.fits FRB20121102 58738.97667 563.8 2.5964 2.84 50 0.0079 0.0224
FRB20121102_0200.fits FRB20121102 58738.97741 566.1 0.7432 5.51 50 0.0093 0.0512
FRB20121102_0201.fits FRB20121102 58738.9778 565.5 0.5270 0.5 100 0.0095 0.0048
FRB20121102_0202.fits FRB20121102 58738.97869 566.4 0.1743 5.32 150 0.01406 0.0748
FRB20121102_0203.fits FRB20121102 58738.98134 567.9 3.5519 3.63 100 0.01496 0.0543
FRB20121102_0204.fits FRB20121102 58738.9872 565.2 1.0840 7.34 150 0.00843 0.0619
FRB20121102_0205.fits FRB20121102 58738.98816 565.5 1.7002 8.47 150 0.00545 0.0462
FRB20121102_0206.fits FRB20121102 58738.9891 566.4 3.8809 2.46 200 0.0073 0.018
FRB20121102_0207.fits FRB20121102 58738.98977 566.8 3.2075 7 50 0.00687 0.0481
FRB20121102_0208.fits FRB20121102 58738.9902 569.3 0.5342 6.05 50 0.00709 0.0429
FRB20121102_0209.fits FRB20121102 58738.99141 568.3 0.1097 8.1 180 0.01024 0.0829
FRB20121102_0210.fits FRB20121102 58738.995 566.7 1.5651 3.69 100 0.00457 0.0169
FRB20121102_0211.fits FRB20121102 58741.87871 565.7 0.1428 2.82 400 0.01089 0.0307
FRB20121102_0212.fits FRB20121102 58741.89634 566.1 4.2045 2.12 300 0.18462 0.3914
FRB20121102_0213.fits FRB20121102 58742.89231 565.6 0.9329 8.39 100 0.00624 0.0524
FRB20121102_0214.fits FRB20121102 58742.90336 564.5 1.0690 4.21 105 0.0044 0.0185
FRB20121102_0215.fits FRB20121102 58743.92509 562.8 3.3248 3.94 100 0.01634 0.0644
FRB20121102_0216.fits FRB20121102 58743.92581 566.5 0.7273 6.2 150 0.00954 0.0592
FRB20121102_0217.fits FRB20121102 58743.9406 566.6 1.4758 3.14 150 0.01919 0.0603
FRB20121102_0218.fits FRB20121102 58743.94263 564.1 1.9594 4.1 400 0.01628 0.0668
FRB20121102_0219.fits FRB20121102 58744.87502 566 3.3659 4.77 140 0.01329 0.0633
FRB20121102_0220.fits FRB20121102 58744.88425 566 2.7010 3.92 120 0.00631 0.0247
FRB20121102_0221.fits FRB20121102 58744.89612 566 0.6872 2.84 80 0.0208 0.0591
FRB20121102_0222.fits FRB20121102 58744.90237 566.3 1.5909 6.58 340 0.00907 0.0597
FRB20121102_0223.fits FRB20121102 58745.93086 567.7 0.5357 4.22 50 0.00805 0.034
FRB20121102_0224.fits FRB20121102 58745.93325 566.4 0.9162 6.4 100 0.00688 0.044
FRB20121102_0225.fits FRB20121102 58745.93452 566.7 1.8966 3.52 100 0.00743 0.0262
FRB20121102_0226.fits FRB20121102 58745.93759 566 2.7633 3.72 100 0.00965 0.0359
FRB20121102_0227.fits FRB20121102 58745.9404 569 4.4028 3.25 100 0.00748 0.0243
FRB20121102_0228.fits FRB20121102 58745.94174 566.2 1.3921 5.21 100 0.00895 0.0466
FRB20121102_0229.fits FRB20121102 58745.94395 568.2 2.7188 9.99 100 0.0052 0.052
FRB20121102_0230.fits FRB20121102 58745.94534 563.4 4.3974 2.66 60 0.02165 0.0576
FRB20121102_0231.fits FRB20121102 58745.94721 567.2 2.9940 2.58 50 0.01108 0.0286
FRB20121102_0232.fits FRB20121102 58745.9476 566.9 4.4542 7.93 100 0.00538 0.0427
FRB20121102_0233.fits FRB20121102 58745.94768 566.2 1.9311 3.35 45 0.00975 0.0327
FRB20121102_0234.fits FRB20121102 58745.95058 564.9 3.9395 5.05 300 0.0076 0.0384
FRB20121102_0235.fits FRB20121102 58745.95123 565.3 2.4797 10.27 200 0.00768 0.0789
FRB20121102_0236.fits FRB20121102 58745.96002 567.5 1.2064 5.68 100 0.01042 0.0592
FRB20121102_0237.fits FRB20121102 58745.96086 566 4.2464 9.55 100 0.00507 0.0484
FRB20121102_0238.fits FRB20121102 58745.96216 565.3 2.4716 4.26 200 0.00936 0.0399
FRB20121102_0239.fits FRB20121102 58745.9653 565 1.8612 3.62 80 0.01498 0.0542
FRB20121102_0240.fits FRB20121102 58746.86415 568.7 4.1699 5.08 100 0.01205 0.0612
FRB20121102_0241.fits FRB20121102 58746.86959 567 2.7811 11.49 400 0.02209 0.2538
FRB20121102_0242.fits FRB20121102 58746.86982 567 2.8310 6.83 400 0.02287 0.1562
FRB20121102_0243.fits FRB20121102 58746.86996 568 2.6960 7.75 50 0.00832 0.0645
FRB20121102_0244.fits FRB20121102 58746.87026 565.5 2.5361 4.96 150 0.00965 0.0479
FRB20121102_0245.fits FRB20121102 58746.87225 568.2 1.2516 3.84 100 0.01288 0.0495
FRB20121102_0246.fits FRB20121102 58746.87421 567.2 1.2864 3.3 100 0.02187 0.0722
FRB20121102_0247.fits FRB20121102 58746.87676 565.4 3.2958 3.75 150 0.01582 0.0593
FRB20121102_0248.fits FRB20121102 58746.88144 565.9 2.8998 5.67 100 0.00553 0.0314
FRB20121102_0249.fits FRB20121102 58746.88301 562.9 2.9836 2.81 300 0.02487 0.0699
FRB20121102_0250.fits FRB20121102 58746.88367 567.8 1.6313 3 34 0.00866 0.026
FRB20121102_0251.fits FRB20121102 58746.88573 566.5 0.7936 4.92 150 0.00758 0.0373
FRB20121102_0252.fits FRB20121102 58746.88867 566.7 2.2947 6.28 400 0.02551 0.1602
FRB20121102_0253.fits FRB20121102 58746.89303 563.6 1.1133 2.95 50 0.0141 0.0416
FRB20121102_0254.fits FRB20121102 58746.89315 566.3 1.7203 6.65 200 0.00297 0.0198
FRB20121102_0255.fits FRB20121102 58746.89415 567.2 4.5206 5.3 100 0.00696 0.0369
FRB20121102_0256.fits FRB20121102 58746.89449 564.7 0.7122 3.99 200 0.0228 0.091
FRB20121102_0257.fits FRB20121102 58746.89593 565.3 1.7717 2.58 300 0.06111 0.1577
FRB20121102_0258.fits FRB20121102 58746.90256 566 4.1600 3.83 210 0.02015 0.0772
FRB20121102_0259.fits FRB20121102 58747.84599 564.9 3.0805 9.59 200 0.00581 0.0557
FRB20121102_0260.fits FRB20121102 58747.8464 570 4.2509 4.8 100 0.00778 0.0373
FRB20121102_0261.fits FRB20121102 58747.8485 565 3.1415 3.52 100 0.02112 0.0743
FRB20121102_0262.fits FRB20121102 58747.85584 565.4 0.8774 7.75 50 0.00664 0.0515
FRB20121102_0263.fits FRB20121102 58747.86076 565.4 1.0834 4.49 200 0.01246 0.056
FRB20121102_0264.fits FRB20121102 58747.86215 565.8 2.8307 2.32 150 0.01034 0.024
FRB20121102_0265.fits FRB20121102 58747.86471 565.6 2.6456 6.74 170 0.01141 0.0769
FRB20121102_0266.fits FRB20121102 58747.8657 564.8 2.0373 4.73 185 0.00681 0.0322
FRB20121102_0267.fits FRB20121102 58747.86672 566.7 0.5945 4.03 150 0.00853 0.0344
FRB20121102_0268.fits FRB20121102 58748.91222 566 3.3601 3.59 150 0.01846 0.0663
FRB20121102_0269.fits FRB20121102 58748.92565 565.3 2.4040 5.36 150 0.00524 0.0281
FRB20121102_0270.fits FRB20121102 58748.92664 567.2 1.7501 6.31 160 0.00661 0.0417
FRB20121102_0271.fits FRB20121102 58748.93031 564.9 3.8434 2.47 100 0.00827 0.0204
FRB20121102_0272.fits FRB20121102 58748.93052 566.3 2.9110 3.2 150 0.00257 0.0082
FRB20121102_0273.fits FRB20121102 58748.93164 569.4 3.1893 4.88 200 0.00843 0.0411
FRB20121102_0274.fits FRB20121102 58748.93784 566.2 4.1185 4.98 120 0.00793 0.0395
FRB20121102_0275.fits FRB20121102 58748.93845 566 0.5550 2.49 110 0.01926 0.0479
FRB20121102_0276.fits FRB20121102 58748.93893 563.7 1.3249 4.47 120 0.00624 0.0279
FRB20121102_0277.fits FRB20121102 58748.93979 566 2.8377 2.4 110 0.04783 0.1147
FRB20121102_0278.fits FRB20121102 58748.94589 564.5 1.9318 9.65 280 0.00891 0.086
FRB20121102_0279.fits FRB20121102 58748.94679 566 1.9071 4.44 180 0.01706 0.0757
FRB20121102_0280.fits FRB20121102 58748.9474 568.5 0.8342 9.55 105 0.0081 0.0774
FRB20121102_0281.fits FRB20121102 58748.94759 565.4 1.8548 5.53 100 0.00379 0.021
FRB20121102_0282.fits FRB20121102 58748.94886 566 3.1757 4.69 90 0.01062 0.0498
FRB20121102_0283.fits FRB20121102 58748.94931 565.4 2.5971 4.8 110 0.01007 0.0484
FRB20121102_0284.fits FRB20121102 58748.94983 567.6 2.5732 7.27 110 0.0109 0.0793
FRB20121102_0285.fits FRB20121102 58748.95086 566 1.2869 3.14 90 0.00816 0.0256
FRB20121102_0286.fits FRB20121102 58749.86271 567.9 2.1985 8.35 100 0.00701 0.0585
FRB20121102_0287.fits FRB20121102 58749.86427 566.9 0.3818 4.72 60 0.00871 0.0411
FRB20121102_0288.fits FRB20121102 58749.86661 567.8 4.2271 2.53 300 0.01515 0.0383
FRB20121102_0289.fits FRB20121102 58749.86779 566.7 4.6022 6.25 120 0.00784 0.049
FRB20121102_0290.fits FRB20121102 58749.86804 565.4 2.7366 3.07 100 0.00683 0.021
FRB20121102_0291.fits FRB20121102 58749.87475 565 3.0190 3.16 200 0.01654 0.0523
FRB20121102_0292.fits FRB20121102 58749.87518 563.6 3.3339 3.37 50 0.00864 0.0291
FRB20121102_0293.fits FRB20121102 58749.87747 566 1.5592 5.22 400 0.01981 0.1034
FRB20121102_0294.fits FRB20121102 58749.87913 567.9 3.2713 7.07 100 0.0086 0.0608
FRB20121102_0295.fits FRB20121102 58749.87979 567.1 2.9559 5.47 100 0.01133 0.062
FRB20121102_0296.fits FRB20121102 58749.88282 566.9 2.3244 6.28 200 0.01118 0.0702
FRB20121102_0297.fits FRB20121102 58749.89254 565.9 2.0921 3.48 140 0.0139 0.0484
FRB20121102_0298.fits FRB20121102 58750.85696 566 4.3124 2.83 260 0.05948 0.1681
FRB20121102_0299.fits FRB20121102 58750.85718 568.5 0.2389 4.6 300 0.01501 0.0691
FRB20121102_0300.fits FRB20121102 58750.8574 568.5 2.4625 4.98 110 0.00583 0.029
FRB20121102_0301.fits FRB20121102 58750.85948 563.7 1.4771 2.64 50 0.01881 0.0497
FRB20121102_0302.fits FRB20121102 58750.86107 564.3 1.0863 5.21 40 0.00993 0.0517
FRB20121102_0303.fits FRB20121102 58750.86208 562.3 3.6439 2.18 220 0.01152 0.0251
FRB20121102_0304.fits FRB20121102 58750.86531 564.6 0.1207 5.54 400 0.02533 0.1403
FRB20121102_0305.fits FRB20121102 58750.86659 566.2 1.0208 8.54 300 0.00772 0.0659
FRB20121102_0306.fits FRB20121102 58750.87267 566.3 4.1655 2.21 130 0.01079 0.0239
FRB20121102_0307.fits FRB20121102 58750.87588 563.3 3.6707 4.87 120 0.00965 0.047
FRB20121102_0308.fits FRB20121102 58750.88589 566.3 0.8042 10.96 180 0.00374 0.041
FRB20121102_0309.fits FRB20121102 58750.88627 569.8 0.6116 8.86 100 0.00576 0.051
FRB20121102_0310.fits FRB20121102 58750.89113 568.8 4.4651 61.56 50 0.01799 1.1075
FRB20121102_0311.fits FRB20121102 58750.89601 566 0.1052 4.02 110 0.00847 0.034
FRB20121102_0312.fits FRB20121102 58750.89622 568.9 2.8886 10.1 170 0.01038 0.1048
FRB20121102_0313.fits FRB20121102 58750.89645 566 0.9421 2.78 280 0.02297 0.064
FRB20121102_0314.fits FRB20121102 58751.88605 565.3 4.3767 2.56 200 0.04741 0.1214
FRB20121102_0315.fits FRB20121102 58751.88985 566.2 1.3331 8.05 50 0.00778 0.0626
FRB20121102_0316.fits FRB20121102 58751.89606 566 2.7540 3.65 110 0.00747 0.0273
FRB20121102_0317.fits FRB20121102 58751.89787 565.3 4.4191 4.98 150 0.01258 0.0627
FRB20121102_0318.fits FRB20121102 58751.90404 566 1.3377 4.5 100 0.01 0.045
FRB20121102_0319.fits FRB20121102 58751.90431 567.2 0.6525 2.43 100 0.00525 0.0128
FRB20121102_0320.fits FRB20121102 58751.90735 566 0.4023 6.15 50 0.00723 0.0445
FRB20121102_0321.fits FRB20121102 58751.90993 561.5 3.2675 7.15 100 0.00493 0.0353
FRB20121102_0322.fits FRB20121102 58751.91079 564.4 4.4307 3.36 100 0.01186 0.0399
FRB20121102_0323.fits FRB20121102 58751.91109 564.4 3.8606 4.41 50 0.0064 0.0282
FRB20121102_0324.fits FRB20121102 58751.91376 564.5 1.0664 3.57 400 0.0117 0.0418
FRB20121102_0325.fits FRB20121102 58751.91945 566.1 2.2207 6.39 130 0.00752 0.0481
FRB20121102_0326.fits FRB20121102 58751.9208 565.1 4.1069 5.94 300 0.00695 0.0413
FRB20121102_0327.fits FRB20121102 58752.85282 564.2 0.4876 4.25 250 0.00961 0.0408
FRB20121102_0328.fits FRB20121102 58752.8536 567.4 1.7484 2.16 300 0.07684 0.166
FRB20121102_0329.fits FRB20121102 58752.85522 555.4 0.6615 7.08 105 0.00308 0.0218
FRB20121102_0330.fits FRB20121102 58752.856 561.5 2.0174 6.57 150 0.00604 0.0396
FRB20121102_0331.fits FRB20121102 58752.85948 565.1 2.1707 7.98 50 0.00731 0.0583
FRB20121102_0332.fits FRB20121102 58752.86034 563.7 4.5954 5.25 220 0.00988 0.0519
FRB20121102_0333.fits FRB20121102 58752.86209 566.6 2.1716 6.27 200 0.00877 0.055
FRB20121102_0334.fits FRB20121102 58752.87052 567.3 2.5578 4.59 250 0.01508 0.0692
FRB20121102_0335.fits FRB20121102 58752.87064 566.5 1.0909 3.5 270 0.05108 0.1783
FRB20121102_0336.fits FRB20121102 58752.87338 567.4 4.4617 4.71 200 0.01145 0.0539
FRB20121102_0337.fits FRB20121102 58752.87364 565.9 1.5629 1.37 100 0.02034 0.0279
FRB20121102_0338.fits FRB20121102 58752.87378 565.7 2.1806 7.16 150 0.00668 0.0478
FRB20121102_0339.fits FRB20121102 58752.87885 566.4 2.1252 3.12 140 0.02886 0.09
FRB20121102_0340.fits FRB20121102 58752.88116 568 0.2969 3.11 450 0.04004 0.1245
FRB20121102_0341.fits FRB20121102 58752.88414 565.2 0.5628 4.48 50 0.00932 0.0418
FRB20121102_0342.fits FRB20121102 58752.88665 564.3 4.4592 3.15 150 0.02319 0.0731
FRB20121102_0343.fits FRB20121102 58752.88687 564.4 1.8776 2.6 220 0.01093 0.0284
FRB20121102_0344.fits FRB20121102 58752.88698 564.3 0.7456 4.6 150 0.01118 0.0514
FRB20121102_0345.fits FRB20121102 58752.88846 563.4 0.0097 9.9 400 0.00692 0.0685
FRB20121102_0346.fits FRB20121102 58753.92768 567.4 1.2670 5.53 70 0.00476 0.0263
FRB20121102_0347.fits FRB20121102 58753.93059 567.1 1.0614 3.92 200 0.01035 0.0406
FRB20121102_0348.fits FRB20121102 58753.93068 567.9 3.6089 3.5 200 0.01502 0.0526
FRB20121102_0349.fits FRB20121102 58753.93376 564.7 0.6172 3.39 160 0.00872 0.0296
FRB20121102_0350.fits FRB20121102 58753.93442 569.3 3.6864 3.28 310 0.0089 0.0292
FRB20121102_0351.fits FRB20121102 58753.93798 567.7 1.6588 3.26 120 0.00725 0.0236
FRB20121102_0352.fits FRB20121102 58753.93928 567.8 1.1772 2.87 100 0.00861 0.0247
FRB20121102_0353.fits FRB20121102 58753.94079 567.1 0.9669 7.44 120 0.0097 0.0722
FRB20121102_0354.fits FRB20121102 58753.94112 562.9 2.3158 4.81 300 0.02063 0.0992
FRB20121102_0355.fits FRB20121102 58753.94543 568 3.9236 2.1 80 0.01031 0.0217
FRB20121102_0356.fits FRB20121102 58753.94629 568.1 0.9342 8.06 110 0.0048 0.0387
FRB20121102_0357.fits FRB20121102 58753.94928 567.6 0.5815 2.28 40 0.01055 0.0241
FRB20121102_0358.fits FRB20121102 58753.95238 565.1 1.3091 7.42 120 0.00947 0.0703
FRB20121102_0359.fits FRB20121102 58754.98428 566 0.4893 3.49 160 0.02719 0.0949
FRB20121102_0360.fits FRB20121102 58754.98708 569.2 1.2337 3.92 400 0.00849 0.0333
FRB20121102_0361.fits FRB20121102 58754.99056 566.7 3.5296 3.59 60 0.01438 0.0516
FRB20121102_0362.fits FRB20121102 58754.9907 568.1 4.2860 3.82 150 0.01289 0.0492
FRB20121102_0363.fits FRB20121102 58754.99448 566.5 2.3784 1.57 360 0.01269 0.0199
FRB20121102_0364.fits FRB20121102 58754.99514 564.9 1.6690 3.13 130 0.0135 0.0423
FRB20121102_0365.fits FRB20121102 58754.99724 566.7 1.6437 3.28 400 0.01878 0.0616
FRB20121102_0366.fits FRB20121102 58754.99752 566.4 0.6105 4.1 150 0.01145 0.0469
FRB20121102_0367.fits FRB20121102 58755.00473 561.9 1.6620 6.03 200 0.00465 0.028
FRB20121102_0368.fits FRB20121102 58755.00658 565.7 4.1140 3.45 100 0.01029 0.0355
FRB20121102_0369.fits FRB20121102 58755.00686 565.2 2.4115 6.22 45 0.00663 0.0412
FRB20121102_0370.fits FRB20121102 58755.01388 568.1 1.1213 2.3 120 0.01525 0.0351
FRB20121102_0371.fits FRB20121102 58755.01393 568.5 0.4012 4.39 300 0.00512 0.0225
FRB20121102_0372.fits FRB20121102 58755.01441 566.3 1.0636 4.05 100 0.00933 0.0378
FRB20121102_0373.fits FRB20121102 58755.01841 564.7 3.8394 2.64 190 0.01624 0.0429
FRB20121102_0374.fits FRB20121102 58755.02051 569.2 4.6207 5.34 200 0.01689 0.0902
FRB20121102_0375.fits FRB20121102 58755.02213 566.2 2.9072 3.55 150 0.02077 0.0737
FRB20121102_0376.fits FRB20121102 58755.0229 564.4 0.5027 3.15 115 0.01611 0.0508
FRB20121102_0377.fits FRB20121102 58755.02339 566 0.2776 3.63 240 0.02006 0.0728
FRB20121102_0378.fits FRB20121102 58756.84317 566 1.1739 2.6 220 0.0254 0.066
FRB20121102_0379.fits FRB20121102 58756.84329 566 2.1476 3.83 170 0.0126 0.0482
FRB20121102_0380.fits FRB20121102 58756.84702 566.7 1.6581 3.4 90 0.01343 0.0457
FRB20121102_0381.fits FRB20121102 58756.84872 565.1 0.5683 3.41 40 0.01185 0.0404
FRB20121102_0382.fits FRB20121102 58756.85209 566.4 1.4503 2.36 60 0.02513 0.0593
FRB20121102_0383.fits FRB20121102 58756.85277 566.2 4.4169 7.03 80 0.01068 0.0751
FRB20121102_0384.fits FRB20121102 58756.85514 566 1.9931 4.9 150 0.01239 0.0607
FRB20121102_0385.fits FRB20121102 58756.85753 566 2.7489 4.06 90 0.00757 0.0307
FRB20121102_0386.fits FRB20121102 58756.85797 565.1 0.7899 2.21 90 0.01265 0.028
FRB20121102_0387.fits FRB20121102 58756.8589 565.7 4.5791 2.37 95 0.01006 0.0238
FRB20121102_0388.fits FRB20121102 58756.85974 564.1 0.9980 5.95 200 0.00393 0.0234
FRB20121102_0389.fits FRB20121102 58756.86064 566.1 0.4693 4.01 350 0.03101 0.1244
FRB20121102_0390.fits FRB20121102 58756.86104 567.2 1.7961 12.1 190 0.00493 0.0596
FRB20121102_0391.fits FRB20121102 58756.86254 566.1 1.5464 3.37 400 0.01625 0.0548
FRB20121102_0392.fits FRB20121102 58756.86532 567.4 3.9016 7.33 190 0.01535 0.1125
FRB20121102_0393.fits FRB20121102 58756.86659 567.3 2.0154 3.59 100 0.01446 0.0519
FRB20121102_0394.fits FRB20121102 58756.8676 567.3 1.5763 3.22 50 0.01418 0.0457
FRB20121102_0395.fits FRB20121102 58756.86956 564.4 0.9883 3.69 140 0.01275 0.0471
FRB20121102_0396.fits FRB20121102 58756.87315 567.3 0.4900 5.81 200 0.02111 0.1227
FRB20121102_0397.fits FRB20121102 58756.87389 565.9 2.8955 9 130 0.00953 0.0858
FRB20121102_0398.fits FRB20121102 58756.87653 566.7 0.2041 5.49 200 0.00829 0.0455
FRB20121102_0399.fits FRB20121102 58756.8777 565.8 3.9097 5.02 230 0.01269 0.0637
FRB20121102_0400.fits FRB20121102 58756.87874 566.7 3.7163 5.31 70 0.00718 0.0381
FRB20121102_0401.fits FRB20121102 58756.88094 566.4 0.8947 3.33 40 0.01468 0.0489
FRB20121102_0402.fits FRB20121102 58757.89884 566 2.5113 6.12 140 0.01427 0.0873
FRB20121102_0403.fits FRB20121102 58757.90141 567 2.7647 6.34 90 0.00453 0.0287
FRB20121102_0404.fits FRB20121102 58757.90386 566.1 1.3238 2.76 180 0.02629 0.0726
FRB20121102_0405.fits FRB20121102 58757.90426 566.1 4.0922 3.74 180 0.01182 0.0442
FRB20121102_0406.fits FRB20121102 58757.90635 565.8 3.1543 2.33 150 0.01611 0.0375
FRB20121102_0407.fits FRB20121102 58757.91478 564.9 3.7430 5.52 120 0.01112 0.0614
FRB20121102_0408.fits FRB20121102 58757.91594 561.2 0.0621 3.98 80 0.00924 0.0368
FRB20121102_0409.fits FRB20121102 58757.91699 567.9 1.1835 5.47 80 0.00807 0.0441
FRB20121102_0410.fits FRB20121102 58757.91916 565.5 3.6007 19.75 80 0.00488 0.0964
FRB20121102_0411.fits FRB20121102 58757.91923 566.9 2.0230 5.95 80 0.00734 0.0437
FRB20121102_0412.fits FRB20121102 58757.91978 564.6 4.2861 4.13 400 0.03347 0.1382
FRB20121102_0413.fits FRB20121102 58757.92414 568 0.9172 5.99 400 0.0196 0.1174
FRB20121102_0414.fits FRB20121102 58757.92447 564.9 1.0551 4.98 150 0.00935 0.0466
FRB20121102_0415.fits FRB20121102 58757.92546 565.4 4.5369 9.2 50 0.00772 0.071
FRB20121102_0416.fits FRB20121102 58757.92926 565.2 4.1266 3.05 450 0.39712 1.2112
FRB20121102_0417.fits FRB20121102 58757.92939 568 1.1817 4.68 150 0.00552 0.0258
FRB20121102_0418.fits FRB20121102 58757.93043 567 0.2854 3.45 300 0.01545 0.0533
FRB20121102_0419.fits FRB20121102 58757.93107 564.8 2.7428 2.37 180 0.00647 0.0153
FRB20121102_0420.fits FRB20121102 58757.9339 564.7 1.5820 2.91 190 0.02252 0.0655
FRB20121102_0421.fits FRB20121102 58757.93409 567.2 3.4956 7.92 115 0.00481 0.0381
FRB20121102_0422.fits FRB20121102 58758.94171 567.4 4.1971 4.55 180 0.01165 0.053
FRB20121102_0423.fits FRB20121102 58758.95195 566.1 2.1474 3.46 100 0.00868 0.03
FRB20121102_0424.fits FRB20121102 58758.9584 565.7 2.7681 1.91 320 0.35048 0.6694
FRB20121102_0425.fits FRB20121102 58758.95904 566.6 0.3917 4.36 100 0.00758 0.0331
FRB20121102_0426.fits FRB20121102 58758.95921 567.6 1.9848 1.8 100 0.00715 0.0129
FRB20121102_0427.fits FRB20121102 58758.96064 567.6 2.9826 9.52 100 0.00741 0.0705
FRB20121102_0428.fits FRB20121102 58758.96135 566.2 4.3479 7.46 100 0.00878 0.0655
FRB20121102_0429.fits FRB20121102 58758.96159 566 0.3154 3.19 115 0.00862 0.0275
FRB20121102_0430.fits FRB20121102 58758.96404 566.2 3.1124 5.15 200 0.01033 0.0532
FRB20121102_0431.fits FRB20121102 58758.96414 569 2.7000 7.33 85 0.00255 0.0187
FRB20121102_0432.fits FRB20121102 58758.96667 565.3 2.5687 4.46 150 0.01102 0.0492
FRB20121102_0433.fits FRB20121102 58758.96702 568 4.4173 6.49 160 0.01055 0.0685
FRB20121102_0434.fits FRB20121102 58758.97225 567.5 2.4748 3.18 100 0.01006 0.032
FRB20121102_0435.fits FRB20121102 58758.9751 564.2 2.0651 3.59 200 0.01506 0.0541
FRB20121102_0436.fits FRB20121102 58758.97645 565.2 2.8060 6.05 280 0.01614 0.0977
FRB20121102_0437.fits FRB20121102 58759.94262 567.1 0.2339 3.8 300 0.01982 0.0753
FRB20121102_0438.fits FRB20121102 58759.94449 564.8 4.2899 5.14 400 0.01306 0.0671
FRB20121102_0439.fits FRB20121102 58759.94491 570 3.6067 3.4 300 0.01185 0.0403
FRB20121102_0440.fits FRB20121102 58759.94825 566.4 0.6175 3.82 300 0.00911 0.0348
FRB20121102_0441.fits FRB20121102 58759.94879 564.3 1.9187 2.44 150 0.02549 0.0622
FRB20121102_0442.fits FRB20121102 58759.95847 567.5 4.1393 4.73 100 0.0086 0.0407
FRB20121102_0443.fits FRB20121102 58759.95866 570 4.5072 7.89 200 0.00551 0.0435
FRB20121102_0444.fits FRB20121102 58759.96085 566.2 3.0032 2.69 100 0.01692 0.0455
FRB20121102_0445.fits FRB20121102 58759.96683 566.5 2.5862 4.92 150 0.01288 0.0634
FRB20121102_0446.fits FRB20121102 58759.96879 568.9 2.8709 8.35 100 0.00664 0.0554
FRB20121102_0447.fits FRB20121102 58759.96943 566 0.6424 3.58 105 0.00801 0.0287
FRB20121102_0448.fits FRB20121102 58759.97445 564.8 3.5696 4.98 110 0.01133 0.0573
FRB20121102_0449.fits FRB20121102 58762.83727 565.8 2.1320 2.62 300 0.01446 0.0379
FRB20121102_0450.fits FRB20121102 58762.84622 564.4 0.6566 2.03 370 0.14109 0.2864
FRB20121102_0451.fits FRB20121102 58762.85632 565.4 1.6781 3.81 250 0.02025 0.0772
FRB20121102_0452.fits FRB20121102 58763.82415 564.7 1.6423 3.16 120 0.01738 0.0549
FRB20121102_0453.fits FRB20121102 58763.84299 565.1 2.1008 3.99 100 0.0091 0.0363
FRB20121102_0454.fits FRB20121102 58766.92944 567.4 3.3591 2.88 130 0.04139 0.1192
FRB20121102_0455.fits FRB20121102 58766.95707 566.3 0.8734 7 400 0.01963 0.1374
FRB20121102_0456.fits FRB20121102 58767.95407 565.1 2.5709 3.5 180 0.0135 0.0446
FRB20121102_0457.fits FRB20121102 58767.96079 564.6 3.2917 3.3 100 0.03825 0.109
FRB20121102_0458.fits FRB20121102 58767.97573 569.8 1.6591 2.85 80 0.00542 0.0572
FRB20121102_0459.fits FRB20121102 58768.91171 569.2 0.9173 3.2 70 0.01609 0.0973
FRB20121102_0460.fits FRB20121102 58768.93974 568 2.8340 3.2 280 0.01921 0.0617
FRB20121102_0461.fits FRB20121102 58772.90647 563.6 4.0108 2.2 110 0.02732 0.0432
FRB20121102_0462.fits FRB20121102 58774.92421 566 2.9677 2.44 105 0.0146 0.0356
FRB20121102_0463.fits FRB20121102 58776.8403 567.4 3.8838 10.71 70 0.00822 0.088
FRB20121102_0464.fits FRB20121102 58776.84098 564.4 1.2192 5.39 70 0.00777 0.0419
FRB20121102_0465.fits FRB20121102 58776.84319 564.4 0.5359 2.12 180 0.04351 0.0922
FRB20121102_0466.fits FRB20121102 58776.855 565.4 2.9935 6.46 70 0.00731 0.0472
FRB20121102_0467.fits FRB20121102 58776.86099 564.7 2.3032 2.9 120 0.00743 0.0216
FRB20121102_0468.fits FRB20121102 58776.87691 565.6 0.8124 2.22 160 0.00933 0.0207
FRB20121102_0469.fits FRB20121102 58776.87755 567.6 3.5502 7.98 75 0.00932 0.0744
FRB20121102_0470.fits FRB20121102 58776.87781 566 0.1444 2.58 300 0.04605 0.119
FRB20180301_0001.fits FRB20180301 59282.46272 517.8 3.3275 12.3 224 0.079 0.9717
FRB20180301_0002.fits FRB20180301 59282.51991 517.81 2.4952 4.1 63 0.044 0.1804
FRB20180301_0003.fits FRB20180301 59282.52654 518 2.2349 1.7 16 0.028 0.0476
FRB20180301_0004.fits FRB20180301 59292.40703 517.83 1.1086 4.1 107 0.105 0.4305
FRB20180301_0005.fits FRB20180301 59292.41184 517.9 4.0180 7.9 142 0.06 0.474
FRB20201124_0001.fits FRB20201124 59482.94673 412.97 2.7003 3.78 150 0.02861 0.108
FRB20201124_0002.fits FRB20201124 59482.95569 412.97 3.7224 3.71 300 0.0307 0.114
FRB20201124_0003.fits FRB20201124 59482.95927 412.97 3.5893 6.77 400 0.22702 1.537
FRB20201124_0004.fits FRB20201124 59482.96238 412.97 2.0929 8.69 250 0.01473 0.128
FRB20201124_0005.fits FRB20201124 59482.96311 412.97 0.8357 6.23 150 0.01011 0.063
FRB20201124_0006.fits FRB20201124 59482.96594 412.97 0.4880 9.44 400 0.19332 1.825
FRB20201124_0007.fits FRB20201124 59482.96783 412.97 3.0348 6.66 150 0.01006 0.067
FRB20201124_0008.fits FRB20201124 59482.98046 412.97 4.7396 9.6 350 0.38298 3.678
FRB20201124_0009.fits FRB20201124 59482.98135 412.97 4.7250 5.51 400 0.033 0.182
FRB20201124_0010.fits FRB20201124 59483.86744 412.18 2.3968 4.14 400 0.03259 0.135
FRB20201124_0011.fits FRB20201124 59483.8701 412.18 0.5039 7.09 450 0.20731 1.47
FRB20201124_0012.fits FRB20201124 59483.87224 412.18 4.4801 5.57 350 0.12014 0.669
FRB20201124_0013.fits FRB20201124 59483.88047 412.18 0.4423 6.46 200 0.01038 0.067
FRB20201124_0014.fits FRB20201124 59483.8957 412.18 2.7318 4.55 100 0.02306 0.105
FRB20201124_0015.fits FRB20201124 59483.89627 412.18 0.0612 8.68 150 0.03019 0.262
FRB20201124_0016.fits FRB20201124 59483.89824 412.18 2.8381 4.04 300 0.05254 0.212
FRB20201124_0017.fits FRB20201124 59483.8993 412.18 3.9064 4.73 250 0.02074 0.098
FRB20201124_0018.fits FRB20201124 59484.81554 411.51 1.9948 5.29 200 0.00812 0.043
FRB20201124_0019.fits FRB20201124 59484.81572 411.51 4.1492 6.33 500 0.10944 0.693
FRB20201124_0020.fits FRB20201124 59484.8162 411.51 1.3943 4.42 200 0.01493 0.066
FRB20201124_0021.fits FRB20201124 59484.81856 411.51 4.8243 6.1 400 0.02854 0.174
FRB20201124_0022.fits FRB20201124 59484.82136 411.51 2.1230 3.59 400 0.22818 0.82
FRB20201124_0023.fits FRB20201124 59484.82186 411.51 0.5458 8.04 200 0.02936 0.236
FRB20201124_0024.fits FRB20201124 59484.82265 411.51 4.0536 7.27 300 0.05555 0.404
FRB20201124_0025.fits FRB20201124 59484.82365 411.51 0.9503 3.06 450 0.0712 0.218
FRB20201124_0026.fits FRB20201124 59484.82496 411.51 3.7418 2.54 200 0.02521 0.064
FRB20201124_0027.fits FRB20201124 59484.82808 411.51 2.8847 2.91 200 0.02368 0.069
FRB20201124_0028.fits FRB20201124 59484.82964 411.51 1.8811 5.92 350 0.08597 0.509
FRB20201124_0029.fits FRB20201124 59484.83066 411.51 0.8569 7.67 100 0.02593 0.199
FRB20201124_0030.fits FRB20201124 59484.83135 411.51 2.3560 2.55 200 0.02704 0.069
FRB20201124_0031.fits FRB20201124 59484.8319 411.51 4.9294 9.63 500 0.35286 3.397
FRB20201124_0032.fits FRB20201124 59484.83209 411.51 1.5283 3.26 200 0.01195 0.039
FRB20201124_0033.fits FRB20201124 59484.83249 411.51 3.9447 7.55 300 0.01484 0.112
FRB20201124_0034.fits FRB20201124 59484.83323 411.51 3.3108 4.48 450 0.33391 1.495
FRB20201124_0035.fits FRB20201124 59484.83476 411.51 0.8373 7.98 300 0.11472 0.916
FRB20201124_0036.fits FRB20201124 59484.83525 411.51 3.3689 10.68 500 0.21182 2.263
FRB20201124_0037.fits FRB20201124 59484.83536 411.51 0.7370 4.42 450 0.68057 3.008
FRB20201124_0038.fits FRB20201124 59484.83653 411.51 4.9067 8.89 250 0.19766 1.757
FRB20201124_0039.fits FRB20201124 59484.83697 411.51 4.4501 3.4 150 0.01939 0.066
FRB20201124_0040.fits FRB20201124 59484.83754 411.51 2.5088 7.05 200 0.01419 0.1
FRB20201124_0041.fits FRB20201124 59484.8399 411.51 0.3685 5.8 350 0.06 0.348
FRB20201124_0042.fits FRB20201124 59484.84035 411.51 0.3013 3.53 200 0.03793 0.134
FRB20201124_0043.fits FRB20201124 59484.8407 411.51 4.8488 4.25 300 0.07298 0.31
FRB20201124_0044.fits FRB20201124 59484.84085 411.51 4.4124 5.08 400 0.02167 0.11
FRB20201124_0045.fits FRB20201124 59484.84309 411.51 4.4885 4.69 300 0.01789 0.084
FRB20201124_0046.fits FRB20201124 59484.84313 411.51 2.4136 6.12 400 0.14354 0.878
FRB20201124_0047.fits FRB20201124 59484.84336 411.51 2.3310 3.87 150 0.01886 0.073
FRB20201124_0048.fits FRB20201124 59484.84419 411.51 2.8879 5.84 400 0.42931 2.506
FRB20201124_0049.fits FRB20201124 59484.84477 411.51 1.7231 4.3 200 0.01024 0.044
FRB20201124_0050.fits FRB20201124 59484.84612 411.51 2.5106 6.99 400 0.30162 2.108
FRB20201124_0051.fits FRB20201124 59484.84633 411.51 1.9113 6.4 100 0.02656 0.17
FRB20201124_0052.fits FRB20201124 59484.84702 411.51 2.9337 5.81 400 0.16846 0.979
FRB20201124_0053.fits FRB20201124 59484.84806 411.51 2.7883 6.74 200 0.02877 0.194
FRB20201124_0054.fits FRB20201124 59484.84908 411.51 0.5312 7.28 100 0.01497 0.109
FRB20201124_0055.fits FRB20201124 59484.84969 411.51 2.2557 3.07 300 0.34016 1.044
FRB20201124_0056.fits FRB20201124 59484.85013 411.51 1.4663 2.88 250 0.57024 1.642
FRB20201124_0057.fits FRB20201124 59484.85061 411.51 3.5011 3.04 400 0.40414 1.229
FRB20201124_0058.fits FRB20201124 59484.85066 411.51 1.8453 13.19 300 0.03708 0.489
FRB20201124_0059.fits FRB20201124 59484.85253 411.51 2.7848 9.86 400 0.29045 2.865
FRB20201124_0060.fits FRB20201124 59485.78533 412.55 1.4799 5.73 200 0.03127 0.179
FRB20201124_0061.fits FRB20201124 59485.78541 412.55 1.9897 4.83 200 0.018 0.087
FRB20201124_0062.fits FRB20201124 59485.78676 412.55 2.7754 7.21 350 0.0334 0.241
FRB20201124_0063.fits FRB20201124 59485.7875 412.55 2.6713 3.3 150 0.02272 0.075
FRB20201124_0064.fits FRB20201124 59485.78764 412.55 1.6791 8.72 350 0.11407 0.995
FRB20201124_0065.fits FRB20201124 59485.78803 412.55 3.0175 4.16 200 0.09997 0.416
FRB20201124_0066.fits FRB20201124 59485.78847 412.55 2.1201 5.36 500 0.08858 0.475
FRB20201124_0067.fits FRB20201124 59485.78854 412.55 2.0278 8.15 500 0.21076 1.718
FRB20201124_0068.fits FRB20201124 59485.78879 412.55 4.1653 4.88 350 0.02993 0.146
FRB20201124_0069.fits FRB20201124 59485.78942 412.55 0.6272 4.54 300 0.07142 0.324
FRB20201124_0070.fits FRB20201124 59485.79069 412.55 1.0980 3.66 300 0.04396 0.161
FRB20201124_0071.fits FRB20201124 59485.79077 412.55 1.2756 8.3 500 0.06737 0.559
FRB20201124_0072.fits FRB20201124 59485.79129 412.55 0.6069 6.08 250 0.03764 0.229
FRB20201124_0073.fits FRB20201124 59485.79462 412.55 4.8612 6.26 250 0.01964 0.123
FRB20201124_0074.fits FRB20201124 59485.79628 412.55 0.4874 9.4 400 0.16216 1.525
FRB20201124_0075.fits FRB20201124 59485.79752 412.55 4.1816 6.62 350 0.1511 1.001
FRB20201124_0076.fits FRB20201124 59485.79845 412.55 1.0572 7.39 300 0.04422 0.327
FRB20201124_0077.fits FRB20201124 59485.79913 412.55 1.8079 3.97 200 0.02089 0.083
FRB20201124_0078.fits FRB20201124 59485.80027 412.55 3.6529 4.83 250 0.0623 0.301
FRB20201124_0079.fits FRB20201124 59485.80083 412.55 0.5508 4.48 350 0.02233 0.1
FRB20201124_0080.fits FRB20201124 59485.80129 412.55 1.5886 4.06 300 0.03007 0.122
FRB20201124_0081.fits FRB20201124 59485.80176 412.55 3.4661 6.45 250 0.16727 1.079
FRB20201124_0082.fits FRB20201124 59485.80197 412.55 2.6063 4 250 0.0885 0.354
FRB20201124_0083.fits FRB20201124 59485.80202 412.55 0.7513 7.78 300 0.04189 0.326
FRB20201124_0084.fits FRB20201124 59485.80252 412.55 4.4643 3.65 400 0.04159 0.152
FRB20201124_0085.fits FRB20201124 59485.8037 412.55 3.9049 7.53 300 0.16536 1.245
FRB20201124_0086.fits FRB20201124 59485.80449 412.55 0.9515 6.53 450 0.08049 0.526
FRB20201124_0087.fits FRB20201124 59485.80463 412.55 0.4211 8.51 300 0.06767 0.576
FRB20201124_0088.fits FRB20201124 59485.80502 412.55 2.0349 7.38 400 0.1015 0.749
FRB20201124_0089.fits FRB20201124 59485.80516 412.55 1.4380 4.07 400 0.20643 0.84
FRB20201124_0090.fits FRB20201124 59485.8055 412.55 4.5371 5.49 500 0.09819 0.539
FRB20201124_0091.fits FRB20201124 59485.80604 412.55 0.4256 2.01 250 0.02288 0.046
FRB20201124_0092.fits FRB20201124 59485.80673 412.55 0.9239 3.27 200 0.02663 0.087
FRB20201124_0093.fits FRB20201124 59485.80764 412.55 3.0362 2.89 200 0.05981 0.173
FRB20201124_0094.fits FRB20201124 59485.8077 412.55 1.6172 6.62 350 0.17298 1.145
FRB20201124_0095.fits FRB20201124 59485.80853 412.55 2.4244 4.99 200 0.01966 0.098
FRB20201124_0096.fits FRB20201124 59485.8086 412.55 2.2274 3.46 200 0.02543 0.088
FRB20201124_0097.fits FRB20201124 59485.8089 412.55 1.9359 7.4 250 0.02379 0.176
FRB20201124_0098.fits FRB20201124 59485.80898 412.55 2.6103 5.46 200 0.03112 0.17
FRB20201124_0099.fits FRB20201124 59485.8093 412.55 4.8104 11.69 450 0.14296 1.671
FRB20201124_0100.fits FRB20201124 59485.80963 412.55 0.9647 4.15 350 0.03327 0.138
FRB20201124_0101.fits FRB20201124 59485.81049 412.55 3.9232 3.86 500 0.24747 0.955
FRB20201124_0102.fits FRB20201124 59485.81246 412.55 0.2918 3.8 400 0.26003 0.989
FRB20201124_0103.fits FRB20201124 59485.81259 412.55 4.9604 5.58 300 0.15388 0.859
FRB20201124_0104.fits FRB20201124 59485.81263 412.55 2.0772 3.9 150 0.02974 0.116
FRB20201124_0105.fits FRB20201124 59485.8128 412.55 3.8942 2.16 250 0.02079 0.045
FRB20201124_0106.fits FRB20201124 59485.81337 412.55 1.9298 3.35 200 0.0421 0.141
FRB20201124_0107.fits FRB20201124 59485.81412 412.55 1.9697 3.69 400 0.02764 0.102
FRB20201124_0108.fits FRB20201124 59485.81436 412.55 3.3692 6.72 450 0.18275 1.228
FRB20201124_0109.fits FRB20201124 59485.81504 412.55 4.2643 2.39 200 0.02221 0.053
FRB20201124_0110.fits FRB20201124 59485.81517 412.55 2.2309 4.54 500 0.53528 2.43
FRB20201124_0111.fits FRB20201124 59485.81546 412.55 1.5814 4.78 150 0.01547 0.074
FRB20201124_0112.fits FRB20201124 59485.8157 412.55 3.4419 8.22 500 0.5477 4.503
FRB20201124_0113.fits FRB20201124 59485.81594 412.55 4.7306 4.03 400 0.07349 0.296
FRB20201124_0114.fits FRB20201124 59485.81643 412.55 1.9233 3 350 0.01764 0.053
FRB20201124_0115.fits FRB20201124 59485.81772 412.55 3.8443 6.21 450 0.09902 0.615
FRB20201124_0116.fits FRB20201124 59485.81813 412.55 0.6751 5.96 150 0.03457 0.206
FRB20201124_0117.fits FRB20201124 59485.8182 412.55 0.3228 2.64 150 0.01629 0.043
FRB20201124_0118.fits FRB20201124 59485.81831 412.55 3.8737 10.91 300 0.05244 0.572
FRB20201124_0119.fits FRB20201124 59485.81878 412.55 4.8637 7.83 200 0.03536 0.277
FRB20201124_0120.fits FRB20201124 59485.82044 412.55 0.8118 5.24 450 0.15476 0.811
FRB20201124_0121.fits FRB20201124 59485.8207 412.55 3.6901 2.67 100 0.02137 0.057
FRB20201124_0122.fits FRB20201124 59485.82105 412.55 2.1136 4.14 250 0.01377 0.057
FRB20201124_0123.fits FRB20201124 59485.82185 412.55 0.5208 2.93 200 0.10873 0.319
FRB20201124_0124.fits FRB20201124 59485.82276 412.55 1.3800 5.01 250 0.03151 0.158
FRB20201124_0125.fits FRB20201124 59485.82381 412.55 1.8504 4.77 400 0.15554 0.742